### v1

# The competitive spectral radius of families of nonexpansive mappings

Abstract.
We consider a new class of repeated zero-sum games in which the payoff is the escape rate of a switched dynamical system, where at every stage, the transition is given by a nonexpansive operator depending on the actions of both players. This generalizes to the two-player (and non-linear) case the notion of joint spectral radius of a family of matrices. We show that the value of this game does exist, and we characterize it in terms of an infinite dimensional non-linear eigenproblem. This provides a two-player analogue of Mañe’s lemma from ergodic control. This also extends to the two-player case results of Kohlberg and Neyman (1981), Karlsson (2001), and Vigeral and the second author (2012), concerning the asymptotic behavior of nonexpansive mappings. We discuss two special cases of this game: order preserving and positively homogeneous self-maps of a cone equipped with Funk’s and Thompson’s metrics, and groups of translations.
Key words and phrases:
Nonexpansive mappings, joint spectral radius, Hilbert and Thompson metrics, zero-sum games, uniform value
1. Introduction
A self map $T$ of a metric space $(X,d)$ is nonexpansive if $d(Tx,Ty)\leqslant d(x,y)$ for all $x,y\in X$. Then, its escape rate is defined by
$\rho(T)=\lim_{k\to\infty}\frac{d(T^{k}\bar{x},\bar{x})}{k}\enspace,$
where $\bar{x}$ is an arbitrary element of $X$. Since $T$ is nonexpansive, the existence of the limit follows from a subadditive argument, and the limit is independent of the choice of $\bar{x}$.
Kohlberg and Neyman proved the following characterization of $\rho(T)$ when $X$ is a star-shaped subset of a normed space.
Theorem 1 ([KN81]).
Let $X$ be a star-shaped subset of a normed space $E$ and $T:X\to X$ be a nonexpansive map, then there exists a continuous linear form $f$ on $E$ of dual norm one such that, for all $x\in X$,
$\rho(T)=\lim_{k\to\infty}\lVert\frac{T^{k}x}{k}\rVert=\inf_{y\in X}\lVert Ty-y\rVert=\lim_{k\to\infty}\langle f,\frac{T^{k}x}{k}\rangle\enspace.$
Moreover, $f$ can be chosen in the weak-* closure of the set of extreme points of the dual unit ball.
Kohlberg and Neyman also show that for all $x\in X$ it is possible to choose $f$ in 1 such that for all $k\in\mathbb{N}^{*}$, we have
$\langle f,T^{k}x\rangle-\langle f,x\rangle\geqslant k\rho(T)\enspace.$
Karlsson extended this result by proving the following Denjoy-Wolff type theorem, which relies on Gromov’s compactification of a metric space by horofunctions [Gro81]. The latter are particular Lipschitz functions of Lipschitz constant $1$ which arise when compactifying the space $X$ as follows: we identify a point $x\in X$ to the function $i(x)\coloneqq d(x,\bar{x})-d(x,\cdot)$ from $X$ to $\mathbb{R}$, where $\bar{x}\in X$ is an arbitrary basepoint; then, we define the horocompactification of $X$ as the closure of $i(X)$ in the topology of pointwise convergence. (This choice of topology was made e.g. in [Rie02, GV12, Kar24]; the topology of uniform convergence on bounded sets was originally considered [Gro81], leading to a different class of horofunctions). The functions of the closure $\overline{i(X)}$ are called metric functionals. The functions of $\overline{i(X)}/i(X)$ are called horofunctions, they constitute the horoboundary.
Theorem 2 (See [Kar01]).
Let $X$ be a proper metric space, $T$ a nonexpansive self-map of $X$, and suppose that every orbit of $T$ in $X$ is unbounded. Then, for all $x\in X$, there is a horofunction $h$ such that
$h(T^{k}x)-h(x)\geqslant k\rho(T)\quad\forall k\geqslant 0\enspace.$
Gaubert and Vigeral refined 2, showing that the horofunction $h$ can be chosen to be independent of $x\in X$ when $X$ satisfies an assumption of non-positive curvature in the sense of Busemann. They consider more generally maps $T$ which are nonexpansive with regard to a hemi-metric of $X$, i.e. a map $d$ taking values in $\mathbb{R}$, verifying the triangular inequality and a weak separation axiom. They established the following maximin characterization of the escape rate.
Theorem 3 (See [GV12]).
Let $T$ be a nonexpansive self-map of a complete metrically star-shaped hemi metric space $(X,d)$. Then
$\rho(T)=\inf_{y\in X}d(Ty,y)=\max_{h\in\mathcal{H}}\inf_{x\in X}(h(Tx)-h(x))$
where $\mathcal{H}$ is the set of metric functionals. Moreover if $d$ is bounded from below and if $\rho(T)>0$ then any function attaining the maximum is a horofunction.
In particular, any horofunction $h$ attaining the above maximum verifies $h(T^{k}x)-h(x)\geqslant k\rho(T)$ for all $k\in\mathbb{N}^{*}$, $x\in X$. Recently, Karlsson further refined this result, showing that when $T$ is an isometry, there exists a metric functional $h$ such that $h(Tx)-h(x)=\rho(T)$ for all $x\in X$, see [Kar24].
In the present paper, we introduce a two-player game extension of the notion of escape rate. We consider a family of nonexpansive self-maps $(T_{ab})_{(a,b)\in\mathcal{A}\times\mathcal{B}}$ of a hemi-metric space $(X,d)$. We call Min and Max the two players. We shall think of $\mathcal{A}$ and $\mathcal{B}$ as the action spaces of players Min and Max, respectively. We assume that Min and Max alternatively select actions $a_{1}\in\mathcal{A},b_{1}\in\mathcal{B},a_{2}\in\mathcal{A},b_{2}\in\mathcal{B},\dots$, after having observed at each stage the actions selected previously by the other player.
The escape rate induced by this infinite sequence of actions is then defined as follows
$\limsup_{k\to\infty}\frac{d(T_{a_{k}b_{k}}\circ\dots\circ T_{a_{1}b_{1}}(\bar{x}),\bar{x})}{k}\enspace,$
where as above, the limit is independent of the choice of $\bar{x}$. Player Max wants to maximize this quantity, while Min wants to minimize it. We call this two-player zero-sum game, Escape Rate Game. We always suppose that the action spaces are compact, and make a technical assumption which is satisfied in particular if the map $(a,b,x)\mapsto T_{ab}(x)$ is continuous (see 7 below). We are interested in the existence of uniform value, a notion studied by Mertens and Neyman [MN81]. Informally, a zero-sum repeated game is said to have a uniform value $\rho$ if both players can guarantee $\rho$ up to an arbitrary precision provided that they play for a sufficiently long time, see 16 below.
To solve this game, we introduce the following Shapley operator $S$, acting on the space $\operatorname{Lip}_{1}$ of Lipschitz functions of constant $1$ from $X$ to $\mathbb{R}$:
$Sv(x)\coloneqq\inf_{a\in\mathcal{A}}\sup_{b\in\mathcal{B}}v(T_{ab}(x))\enspace.$
Our main result is the following maximin characterization of the value of the escape rate game. Again, $\bar{x}$ denotes an arbitrary point in $X$.
Theorem 4.
The escape Rate Game has a uniform value $\rho$ given by
$\rho=\lim_{k\to\infty}\frac{[S^{k}d(\cdot,\bar{x})](\bar{x})}{k}=\max\{\lambda\in\mathbb{R}\mid\exists\;v\in\operatorname{Lip}_{1},\lambda+v\leqslant Sv\}=\max_{v\in\operatorname{Lip}_{1}}\inf_{x\in X}(Sv(x)-v(x))\enspace.$
We show in particular that Player Max has an optimal stationary strategy, and that Player Min has an optimal (non stationary) strategy.
We say that a function $f\in\operatorname{Lip}_{1}$ is distance-like if there exists a constant $\alpha$ such that $f(x)\geqslant d(x,\bar{x})+\alpha$. We denote by $\mathscr{D}$ the set of distance like functions. We say that a nonexpansive self map $T$ of a hemi metric space $(X,d)$ an almost-isometry with constant $\gamma\in\mathbb{R}$ if it satisfies $d(T(x),T(y))\geqslant d(x,y)-\gamma$, for all $x,y\in X$.
Theorem 5.
Suppose that the maps $T_{ab}$ constitute a family of almost-isometries with a uniform constant. Then. the value of the escape rate game satisfies
$\rho=\inf\{\lambda\in\mathbb{R}\mid v\in\mathscr{D},\,\lambda+v\geqslant Sv\}=\inf_{v\in\mathscr{D}}\sup_{x\in X}(Sv(x)-v(x))\enspace.$
The infimum is not achieved in general.
The introduction of escape rate games is motivated by matrix multiplication games, defined by Asarin et al. in [ACD+16]. In these games, one considers two compact sets of matrices $\mathcal{A}$ and $\mathcal{B}$ with entries in $\mathbb{C}$, and the players Min and Max alternatively select matrices $a_{1}\in\mathcal{A}$, $b_{1}\in\mathcal{B}$, $a_{2}\in\mathcal{A}$, …Player Max wants to maximize the payoff
(1) $\limsup_{k}\|a_{1}b_{1}\dots a_{k}b_{k}\|^{1/k}$
[…]
If $\mathcal{A}$ and $\mathcal{B}$ are included in $\operatorname{GL}_{n}$, we associate to a pair of matrices $(a,b)\in\mathcal{A}\times\mathcal{B}$ the self map of the open cone of Hermitean positive definite matrices $H_{n}^{++}$ defined by $T_{ab}(x)=b^{*}a^{*}xab$. Moreover, we equip $H_{n}^{++}$ with the Funk metric [PT08], defined by
$\operatorname{Funk}(x,y)=\log\inf\{\alpha>0\mid y\leqslant\alpha x\}=\log\lambda_{\max}(xy^{-1})\enspace,$
[…]
$\log\|a_{1}b_{1}\dots a_{k}b_{k}\|=\frac{1}{2}\operatorname{Funk}(T_{a_{k}b_{k}}\circ\dots\circ T_{a_{1}b_{1}}(I),I)$
where $I$ is the identity matrix. Therefore, matrix multiplications games in $\operatorname{GL}_{n}$ reduce to escape rate games. By applying 4 and 5, we obtain the existence together with characterizations of the uniform value of matrix multiplication games.
In the special case in which Player Min is a dummy, i.e. $\mathcal{A}=\{I\}$, the value of the escape rate game coincides with the joint spectral radius introduced by Rota and Strang [RS60]. When Player Max is a dummy, it coincides with the notion of lower spectral radius or joint spectral subradius introduced much later by [Gur95]. Therefore, we call competitive spectral radius the value of an escape rate game, as it generalizes both notions. Our approach allows us more generally to define the competitive spectral radius of a family of non-linear maps, positively homogeneous of degree one, acting on an open finite dimensional cone, and preserving the order of this cone. Moreover, if the maps extend continuously to the closure of the cone and $(a,b,x)\mapsto T_{ab}(x)$ is continuous, we show that the dual characterization 5 holds without a “quasi-isometry” assumption. We obtain in particular the existence of the value for families of non necessarily invertible nonnegative matrices.
The notion of joint spectral radius of families of isometries introduced recently by Breuillard and Fujiwara [BF21], as well as the notions of joint spectral radius of family of topical maps defined by Bousch and Mairesse [BM01], are also special (one-player) cases of competitive spectral radii.
Theorem 5 may be thought of as a two player version of the Mañé-Conze-Guivarch lemma arising in ergodic control, see [Mañ92] and the discussion by Bousch in [Bou11]. It is also closely related with the weak-KAM solutions defined by Fathi [Fat97].
We illustrate these results by solving a “simple” special class of escape rate game, in which the nonexpansive dynamics consists of translations acting on the Euclidean space. We give an explicit formula for a maximizing function $v$ in the maximin characterization of 5, which is based on a result of discrete geometry (Shapley-Folkman lemma).
Concrete examples of competitive spectral radii arise in the study of topological entropies of transition systems from program and systems verification [ACD+16] and population dynamics, in particular in problems from mathematical biology and epidemiology, see in particular [BCF+14, CGG14, CWW+08]. They also arise in robust Markov decision processes, in which the second player represents the uncertainty, see [GGC23].
2. The Competitive Spectral Radius
2.1. Hemi-metric spaces
We shall need the following notion of asymmetric metric [GV12].
Definition 6 (Hemi-metric).
A map $d:X\times X\to\mathbb{R}$ is called a hemi-metric on the set $X$ if it satisfies the two following conditions:
  1. (1)
    $\forall(x,y,z)\in X^{3},\;d(x,z)\leqslant d(x,y)+d(y,z)$ (triangular inequality)
  2. (2)
    $\forall(x,y)\in X^{2},\;d(x,y)=d(y,x)=0$ if and only if $x=y$
Other definitions requiring for $d$ to be nonnegative have appeared in the literature (see [PT08] for instance), but this is undesirable in our approach. A basic example of hemi-metric on $\mathbb{R}^{n}$ is the map $\delta_{n}(x,y):=\max_{i\in[n]}(x_{i}-y_{i})$.
Given a hemi-metric $d$, the map
$d^{\circ}(x,y)=\max(d(x,y),d(y,x))\enspace,$
is always a metric on $X$. We will refer to it as the symmetrized metric of $d$. In the sequel, we equip $X$ with the topology induced by $d^{\circ}$.
Given two hemi-metric spaces $(X,d)$ and $(Y,d^{\prime})$, we introduce the notion of a $1$-Lipschitz or nonexpansive function $v$ with respect to $d$ and $d^{\prime}$. Such a function verifies $d^{\prime}(v(x),v(y))\leqslant d(x,y)$ for all $x,y\in X$. We will especially be interested in the nonexpansive functions from $(X,d)$ to $(\mathbb{R},\delta_{1})$.
It is easy to verify that a Lipschitz function between two hemi-metric spaces is Lipschitz with respect to the symmetrized metrics generated by their respective hemi-metrics. In particular, such a function is continuous with respect to the topologies induced by these symmmetrized metrics.
2.2. The Escape Rate Game
The Escape Rate Game is the following two-player deterministic perfect information game.
We fix two non-empty compact sets $\mathcal{A}$ and $\mathcal{B}$, that will represent the action spaces of the two players. We also fix a hemi-metric space $(X,d)$, which will play the role of the state space of the game. To each pair $(a,b)\in\mathcal{A}\times\mathcal{B}$, we associate a nonexpansive self-map $T_{ab}$ of $(X,d)$. The initial state $x_{0}\in X$ is given. We construct inductively a sequence of states $(x_{k})_{k\geqslant 0}$ as follows. Both players observe the current state $x_{k}$. At each turn $k\geqslant 1$, player Min chooses an action $a_{k}\in\mathcal{A}$ and then, after having observed the action $a_{k}$, player Max chooses an action $b_{k}\in\mathcal{B}$. The next state is given by $x_{k}=T_{a_{k}b_{k}}(x_{k-1})$.
First, we define the game in horizon $k$, denoted by $\Gamma_{k}$. There are $k$ successive turns, and player Min pays to Max the following amount:
(2) $J_{k}(a_{1}b_{1}\dots a_{k}b_{k})=d(T_{a_{k}b_{k}}\circ\dots\circ T_{a_{1}b_{1}}(x_{0}),x_{0})\enspace.$
Player Max wishes to maximize this quantity, whereas player Min wishes to minimize it.
We also define the escape rate game, denoted $\Gamma_{\infty}$. There, an infinite number of turns are played, and player Min wishes to minimize the escape rate
$J_{\infty}(a_{1}b_{1}a_{2}b_{2}\dots)\coloneqq\limsup_{k\to\infty}\frac{J_{k}(a_{1}b_{1}\dots a_{k}b_{k})}{k}\enspace.$
while player Max wants to maximize it. It follows from the nonexpansiveness of the maps $T_{ab}$ that this limit is independent of the choice of $x_{0}$.
A strategy is a map that assigns a move to any finite history of states and actions (for a sequence of turns). To formalize this notion, it will be convenient to use the notation $\mathcal{C}^{*}\coloneqq\cup_{k\geqslant 0}\mathcal{C}^{k}$, where $\mathcal{C}^{k}=\mathcal{C}\times\dots\times\mathcal{C}$, for the set of finite sequences of elements of $\mathcal{C}$.
Then, a strategy for Min is a map
$\sigma:(X\times\mathcal{A}\times\mathcal{B})^{*}\times X\to\mathcal{A},$
where the action chosen at turn $k+1$ depends on the history $(x_{0},a_{1},b_{1},x_{1},\ldots,a_{k},b_{k},x_{k})$ of states and actions. Similarly, a strategy for Max is a map
$\tau:(X\times\mathcal{A}\times\mathcal{B})^{*}\times X\times\mathcal{A}\to\mathcal{B}\enspace.$
The set of strategies of Min (resp Max) will be written $\mathcal{A}^{s}$ (resp $\mathcal{B}^{s}$). A strategy $\sigma\in\mathcal{A}^{s}$ of Min is called stationary if it is independent of the turn $k$ and depends only on the position in the space, i.e. if for all $x\in X$, the restriction of $\sigma$, $\sigma_{x}:(X\times\mathcal{A}\times\mathcal{B})^{*}\times\{x\}\to\mathcal{A}$ is constant, it is then positional, since the action chosen at turn $k+1$ only depends on the position of $x_{k}$. Similarly, a strategy of Max is stationary if it is independent of the turn $k$, and depends only on the position in the space and the last play of Min. In general, since the game is deterministic, any strategy can be rewritten as a function of the initial state and of the history of actions, but then the stationarity may not be visible.
For every pair $(\sigma,\tau)\in\mathcal{A}^{s}\times\mathcal{B}^{s}$, we denote by $J_{k}(\sigma,\tau)$ the payoff of $\Gamma_{k}$, as per (2), assuming that the sequence of actions $a_{1},b_{1},a_{2},b_{2},\dots$ is generated according to the strategies $\sigma$ and $\tau$, and similarly, we denote by $J_{\infty}(\sigma,\tau)$ the escape rate defined above.
Recall that the value of a game is the unique quantity that the players can both guarantee. Formally a game with strategy spaces $\mathcal{A}^{s}$ and $\mathcal{B}^{s}$ and payoff function $\mathcal{A}^{s}\times\mathcal{B}^{s}\to\mathbb{R}$, $(\sigma,\tau)\mapsto J(\sigma,\tau)$, has a value $\lambda\in\mathbb{R}$ if $\forall\epsilon>0,\;\exists\sigma^{*}\in\mathcal{A}^{s},\tau^{*}\in\mathcal{B}^{s}$ such that
$J(\sigma^{*},\tau)\leqslant\lambda+\epsilon\text{ and }J(\sigma,\tau^{*})\geqslant\lambda-\epsilon$
for all strategies $\sigma\in\mathcal{A}^{s}$ and $\tau\in\mathcal{B}^{s}$. Such strategies are called $\epsilon$-optimal strategies. When $\epsilon=0$, one gets optimal strategies, and in particular, $\lambda=J(\sigma^{*},\tau^{*})$.
2.3. Examples
We now describe several examples of escape rate games.
2.3.1. Order preserving positively homogeneous self-maps of cones
Let $C\subset\mathbb{R}^{n}$ be a closed convex pointed cone of non-empty interior. We define the partial order $x\leqslant_{C}y$ if $y-x\in C$. We equip the interior of $C$, $\operatorname{int}C$, with the Funk metric
$d(y,z)=\operatorname{Funk}(y,z)\coloneqq\log\inf\{\alpha>0\mid y\leqslant\alpha z\}\enspace.$
Notice that we could define this hemi-metric on $C\times\operatorname{int}C$ however there would be no clear topology on $C$ as the symmetrized metric would not be defined for all pair of points. The symmetrized metric is known as Thompson’s part metric:
$d^{\circ}(y,z)=\operatorname{Th}(y,z):=\max(\operatorname{Funk}(y,z),\operatorname{Funk}(z,y))\enspace.$
A related “metric” of interest is Hilbert’s projective metric:
$\operatorname{Hil}(y,z):=\frac{\operatorname{Funk}(y,z)+\operatorname{Funk}(z,y)}{2}\enspace.$
It is a projective metric, meaning that $\operatorname{Hil}(y,z)=0$ iff $y$ and $z$ are proportional.
For instance, if $C=(\mathbb{R}_{\geqslant 0})^{n}$, then $\operatorname{int}C=(\mathbb{R}_{>0})^{n}$, and for all $y,z\in(\mathbb{R}_{>0})^{n}$, setting $\operatorname{top}(x):=\max_{i}x_{i}$,
$\operatorname{Funk}(y,z)=\log\max_{i\in[n]}\frac{y_{i}}{z_{i}}=\operatorname{top}(\log y-\log z),\qquad\operatorname{Th}(y,z)=\|\log y-\log z\|_{\infty}\enspace.$
If $C=\mathcal{H}_{n}$ is the cone of Hermitean positive semidefinite matrices, then, $\operatorname{int}C$ is the cone of positive definite Hermitean matrices, and
$\operatorname{Funk}(y,z)=\log\lambda_{\max}(yz^{-1})\enspace.$
The Funk metric arises in Hilbert’s geometry, see [PT08].
Remark 1.
Note that $\operatorname{Funk}(y,z)$ may take negative values. In contrast, Papadopoulos and Troyanov require weak metrics to be nonnegative, and so define the Funk metric to be $\max(0,\operatorname{Funk}(y,z))$. We prefer to work with the unsigned version as it captures more information.
We call order preserving a self-map $f$ of $\operatorname{int}C$ such that $x\leqslant_{C}y\implies f(x)\leqslant_{C}f(y)$. It is called positively homogeneous (of degree one) if for all $\lambda>0,x\in C$, $f(\lambda x)=\lambda f(x)$. Satisfying both these conditions is equivalent to the map being nonexpansive regarding the Funk metric. In particular, we denote by $\operatorname{End}(C)$ the set of linear self-maps of $\mathbb{R}^{n}$ which preserve the open cone $\operatorname{int}C$. Any element of $\operatorname{End}(C)$ is nonexpansive for the Funk metric.
For instance, if $C=(\mathbb{R}_{\geqslant 0})^{n}$, then, $\operatorname{End}(C)=\mathcal{M}^{n}_{+}$, the set of nonnegative matrices $M$ with at least one non-zero entry per column, acting on row vectors by $x\mapsto xM$. If $C=\mathcal{H}_{n}^{+}$, the set $\operatorname{End}(C)$ is harder to describe. It contains all completely positive maps of the form
$T(x)=\sum_{i=1}^{k}a_{i}^{*}xa_{i}$
where $a_{1},\dots,a_{k}\in\mathbb{C}^{n\times n}$ are such that $\sum_{i=1}^{k}a_{i}^{*}a_{i}$ is invertible.
2.3.2. Families of Nonnegative Matrices
The special case $C=(\mathbb{R}_{\geqslant 0})^{n}$ is of interest. We consider two non-empty compact subsets $\mathcal{A}$ and $\mathcal{B}$ of +$\mathcal{M}^{n}_{+}$. Each pair of actions $(A,B)\in\mathcal{A}\times\mathcal{B}$ induces the self-map of $\mathbb{R}_{>0}^{n}$ given by
$T_{AB}(x)\coloneqq xAB\enspace,$
in which $x$ is interpreted as a row vector. Since $T_{AB}\in\operatorname{End}((\mathbb{R}_{\geqslant 0})^{n})$, it is nonexpansive in the Funk hemi-metric.
This yields an escape rate game, in which the payoff is given by
$\limsup_{k\to\infty}\frac{\operatorname{Funk}(x_{k},x_{0})}{k}$ $=\limsup_{k\to\infty}\log\max_{i\in\llbracket 1,n\rrbracket}(\frac{(x_{k})_{i}}{(x_{0})_{i}})^{\frac{1}{k}}$
$=\log(\limsup_{k\to\infty}(\max_{i\in\llbracket 1,n\rrbracket}(x_{k})_{i})^{1/k})$
$=\log(\limsup_{k\to\infty}\lVert x_{k}\rVert^{1/k}_{\infty})\enspace,$
where
$x_{k}=x_{0}A_{1}B_{1}\dots A_{k}B_{k}\enspace.$
Taking $x_{0}=(1,\dots,1)$ we get $\lVert x_{k}\rVert_{\infty}=\lVert A_{1}B_{1}\dots A_{k}B_{k}\rVert$, where for $A=(a_{ij})\in\mathbb{R}^{n\times n}$, $\lVert A\rVert=\max_{i}\sum_{k=1}^{n}\lvert a_{ik}\rvert$, and we recover the payoff of a matrix multiplication game.
The reverse Funk metric
$\operatorname{RFunk}(x,y):=\operatorname{Funk}(y,x)$
is also of interest. It yields an escape rate game in which the payoff is given by
$\limsup_{k\to\infty}\frac{\operatorname{RFunk}(x_{k},x_{0})}{k}$ $=\limsup_{k\to\infty}\log\max_{i\in\llbracket 1,n\rrbracket}(\frac{(x_{0})_{i}}{(x_{k})_{i}})^{1/l}$
$=-\liminf_{k\infty}\log\min_{i\in\llbracket 1,n\rrbracket}(\frac{(x_{k})_{i}}{(x_{0})_{i}})^{1/l}\enspace.$
Then, one player wants to maximize the smallest growth rate of the coordinates of $x^{k}$, whereas the other player wants to maximize it.
Such models arise in population dynamics, control of growth processes, and epidemiology: $x_{k}$ represents a population profile at time $k$; one player wishes to minimize the growth rate of the population whereas the other player wishes to maximize it, see [BCF+14, CGG14, CWW+08].
2.3.3. Families of Matrices in $\operatorname{GL}_{n}(\mathbb{C})$
Every matrix $M\in\operatorname{GL}_{n}(\mathbb{C})$ yields a congruence operator,
$\phi_{M}:X\mapsto M^{*}XM$
which is an automorphism of the cone of positive semidefinite Hermitean matrices, i.e., $\phi_{M}\in\operatorname{Aut}(\mathcal{H}_{n}^{+})$.
We say that a function $\nu$ on $\mathbb{R}^{n}$ is a symmetric Gauge function if it is invariant by permutation of the variables, convex and positively homogeneous of degree $1$.
Every symmetric Gauge function $\nu$ yields a hemi-metric of Finsler type on $\operatorname{int}(\mathcal{H}_{n}^{+}d)$ ([LN12])
$d_{\nu}(X,Y)=\nu(\log(\operatorname{Spec}(XY^{-1})))\enspace.$
If $M\in\operatorname{GL}_{n}(\mathbb{C})$, then $\phi_{M}$ is an isometry of $(\operatorname{int}(\mathcal{H}_{n}^{+}d),d_{\nu})$. Indeed for all $X,Y\in\operatorname{int}(\mathcal{H}_{n}^{+}d)$
$d_{\nu}(\phi_{M}(x),\phi_{M}(Y))$ $=\nu(\log(\operatorname{Spec}((M^{*}XM)(M^{*}YM)^{-1})))$
$=\nu(\log(\operatorname{Spec}(M^{*}XMM^{-1}Y^{-1}(M^{*})^{-1})))$
$=\nu(\log(\operatorname{Spec}(M^{*}XY^{-1}(M^{*})^{-1})))=d_{\nu}(X,Y)$
Consider $\mathcal{A},\mathcal{B}\subset\operatorname{GL}_{n}$ and the operators $(\phi_{AB})_{(A,B)\in\mathcal{A}\times\mathcal{B}}$, which are isometries, as $\phi_{AB}=\phi_{B}\circ\phi_{A}$ We denote by $\operatorname{Sing}$ the set of singular values of a matrix. We get that
$d_{\nu}(\phi_{A_{k}B_{A}}\circ\dots\circ\phi_{A_{1}B_{1}}(I),I)=2\nu(\log(\operatorname{Sing}(A_{1}B_{1}\dots A_{k}B_{k}))$
Therefore, depending on the Gauge function we can get some insights about the spectrum and singular values of the infinite product of matrices. In particular, when $\nu(x)=\operatorname{top}(x)=\max_{i}x_{i}$,
$d_{\operatorname{top}}(\phi_{A_{k}B_{A}}\circ\dots\circ\phi_{A_{1}B_{1}}(I),I)=2\log\|A_{1}B_{1}\dots A_{k}B_{k}\|$
[…]
$\limsup_{k\to\infty}\lVert A_{1}B_{1}\dots A_{k}B_{k}\rVert^{\frac{1}{k}}$
whereas the other wants to minimize it.
2.3.4. Vector Addition Games
[…]
$J_{\infty}(a_{1}b_{1}a_{2}b_{2}\dots)=\limsup_{k\to\infty}\frac{\|a_{1}+b_{2}+\dots+a_{k}+b_{k}\|}{k}\enspace.$
3. Main Results
To prove the existence and characterize the value of escape rate games, we use an operator approach. We make the following assumption.
Assumption 7.
For all $x\in X$, the maps $b\mapsto T_{ab}(x)$ and $a\mapsto T_{ab}(x)$ are continuous, and for all compact sets $K$, the set $\{T_{ab}(x)\ |\ (a,b,x)\in\mathcal{A}\times\mathcal{B}\times K\}$ is compact.
This is trivially verified if $(a,b,x)\mapsto T_{ab}(x)$ is continuous.
We denote by $\operatorname{Lip}_{1}$ the set of $1$-Lipschitz functions from the hemi-metric space $(X,d)$ to $(\mathbb{R},\delta_{1})$ (recall that $\delta_{1}(x,y)=x-y$). We define the following operator $S$ acting on $\operatorname{Lip}_{1}$,
(3) $Sv(x)\coloneqq\inf_{a\in\mathcal{A}}\sup_{b\in\mathcal{B}}v(T_{ab}(x))$
As proven in 9 below, 7 ensures that $S$ preserves the set $\operatorname{Lip}_{1}$ and is continuous for the topology of uniform convergence on compact sets.
The main goal of this section is to prove a Collatz-Wielandt-type theorem. Indeed, we want to prove that the game has a value and that it is equal to
(4) $\max\{\lambda\in\mathbb{R}\mid\exists v\in\operatorname{Lip}_{1},\;\lambda+v\leqslant Sv\}\enspace.$
3.1. Preliminary Results
In this subsection, we show a few intermediary results.
Proposition 8.
Let $x_{0}\in X$, $\mathcal{L}_{x_{0}}$ the subset of $\operatorname{Lip}_{1}$ of functions equal to 00 in $x_{0}$ is compact for the topology of compact convergence (otherwise called uniform convergence on compact sets).
Proof.
This set is obviously equi-continuous since it is equi-Lipschitz (for the hemi-metrics as well as the symmetrized metrics). Furthermore $\forall x\in X,\forall v\in\mathcal{L}_{x_{0}}$, we have $|v(x)|=|v(x)-v(x_{0})|\leqslant d^{\circ}(x,x_{0})$. So $\forall x$, $\{v(x)\mid v\in\mathcal{L}_{x_{0}}\}$ is a bounded set in $\mathbb{R}$ and therefore it is relatively compact. Now, by the general Ascoli-Arzela theorem (for the statement and its proof see [Wan21] for instance) we know that $\mathcal{L}_{x_{0}}$ is relatively compact for the topology of compact convergence (which coincides with the topology of pointwise convergence as $\mathcal{L}_{x_{0}}$ is equi-continuous [Kel75, Th 15 on p.232]). It is closed for this topology so it is compact. ∎
Proposition 9.
Under 7, the operator $S$ preserves $\operatorname{Lip}_{1}$ and is continuous from $\operatorname{Lip}_{1}$ to itself for the topology of compact convergence.
Proof.
By 7, for any $x\in X$, and $a\in\mathcal{A}$, the map $b\in\mathcal{B}\mapsto T_{ab}(x)\in X$ is continuous. Let $v\in\operatorname{Lip}_{1}$. Since then $v$ is continuous from $X$ to $\mathbb{R}$ and $\mathcal{B}$ is compact, we get that $Sv(x)\leqslant\sup_{b\in\mathcal{B}}v(T_{ab}(x))<+\infty$. Similarly, $Sv(x)>-\infty$. This implies that $Sv(x)\in\mathbb{R}$ for all $x\in X$. By assumption the maps $T_{ab}$ are nonexpansive, hence all maps $x\mapsto v(T_{ab}(x))$ are in $\operatorname{Lip}_{1}$, so that $Sv$ also belongs to $\operatorname{Lip}_{1}$, as $\operatorname{Lip}_{1}$ is preserved by suprema and infima.
Let us now show that $S$ is continuous for the topology of compact convergence.
Let $I$ be a directed set and consider a net $(v_{i})_{i\in I}$ converging to a function $v$ that belongs to $\operatorname{Lip}_{1}$. Let $K$ be a compact subset of $X$. We have
(5) $\sup_{x\in K}|Sv_{i}(x)-Sv(x)|$ $\leqslant\sup_{(x,a,b)\in K\times\mathcal{A}\times\mathcal{B}}|v_{i}(T_{ab}(x))-v(T_{ab}(x))|$
(6) $=\sup_{y\in Y}|v_{i}(y)-v(y)|$
where, by 7, $Y:=\{T_{ab}(x)|(x,a,b)\in K\times\mathcal{A}\times\mathcal{B}\}$ is compact. Therefore the RHS converges to 0 and so does the LHS. So $Sv_{i}\to Sv$ in the topology of compact convergence. ∎
Proposition 10.
There exists at least an additive eigenvalue in $\operatorname{Lip}_{1}$ for $S$, i.e
$\exists\lambda\in\mathbb{R}\text{ such that }\exists v\in\operatorname{Lip}_{1},Sv=\lambda+v$
Proof.
Let us remind that for a topological space $X$, the space of continuous real-valued functions, $\mathcal{C}(X)$, equipped with the topology of uniform convergence on compact sets, which is defined by the family of semi-norms $\phi_{K}:f\mapsto\sup\{\lvert f(x)\rvert\mid x\in K\}$ (where $K$ varies over the directed set of all compact subsets of $X$), is a locally convex topological vector space. Let us define $\tilde{S}v\coloneqq Sv-Sv(x_{0})$. It sends continuously $\mathcal{L}_{x_{0}}$ into itself. As $\mathcal{L}_{x_{0}}$, is a convex and compact subset of $\mathcal{C}(X)$, we can conclude that $\tilde{S}$ has a fixed point in $\mathcal{L}_{x_{0}}$ thanks to the generalization of the Schauder fixed point theorem by Tychonoff [Tyc35]. ∎
Notice that this result becomes trivial if $d$ is a nonnegative hemi metric since the null function would be an additive eigenvector with respect to the eigenvalue 00.
By the triangular inequality, the map $d(\cdot,x_{0}):x\mapsto d(x,x_{0})$ is in $\operatorname{Lip}_{1}$. We define
$s_{k}\coloneqq[S^{k}d(\cdot,x_{0})](x_{0})\enspace.$
Proposition 11.
We have
(7) $\rho:=\inf_{k\geqslant 1}\frac{s_{k}}{k}=\lim_{k}\frac{s_{k}}{k}\in\mathbb{R}\enspace.$
Proof.
We first show the following subadditivity property:
(8) $s_{k+l}\leqslant s_{k}+s_{l},\text{ for all }k,l\geqslant 1\enspace.$
Indeed, we have $S^{l}d(\cdot,x_{0}))\leqslant[S^{l}d(\cdot,x_{0})](x_{0})+d(\cdot,x_{0})=s_{l}+d(\cdot,x_{0})$ since $S^{l}d(\cdot,x_{0})$ is $1$-Lipschitz. Then, since $S^{k}$ is order preserving, and commutes with the addition of a constant, $S^{k+l}d(\cdot,x_{0})=S^{k}(S^{l}d(\cdot,x_{0}))\leqslant S^{k}(s_{l}+d(\cdot,x_{0}))=s_{l}+S^{k}(d(\cdot,x_{0}))$, and so, $[S^{k+l}d(\cdot,x_{0})](x_{0})\leqslant s_{l}+s_{k}$, showing (8). Then, it follows from Fekete subadditive lemma that the limit $\lim_{k}\frac{s_{k}}{k}$ does exists and that it coincides with $\rho:=\inf_{k}\frac{s_{k}}{k}$. In particular $\rho\leqslant s_{1}<+\infty$.
Let $v=-d(x_{0},\cdot)$. By the triangular inequality, $v$ is in $\operatorname{Lip}_{1}$, so that $Sv(x_{0})\in\mathbb{R}$. We set
$\alpha:=[S(-d(x_{0},\cdot)](x_{0})=\inf_{a}\sup_{b}-d(x_{0},T_{ab}(x_{0}))\in\mathbb{R}$
Again by the triangular inequality, we have $0=d(x_{0},x_{0})\leqslant d(x_{0},\cdot)+d(\cdot,x_{0})$.
We deduce that
$s_{1}=[Sd(\cdot,x_{0})](x_{0})=\inf_{a}\sup_{b}d(T_{ab}(x_{0}),x_{0})\geqslant\alpha\enspace.$
We now show, by induction on $k$, that $s_{k}\geqslant\alpha k$ for all $k$. Since $S$ preserves $\operatorname{Lip}_{1}$, we have, for every $k\in\mathbb{N}^{*}$, $S^{k}d(\cdot,x_{0})\in\operatorname{Lip}_{1}$, and so
$[S^{k}d(\cdot,x_{0})](\cdot)\geqslant[S^{k}d(\cdot,x_{0})](x_{0})-d(x_{0},\cdot)=s_{k}-d(x_{0},\cdot)$
So, as $S$ is order preserving and commutes with the addition of a constant, we have
$[S^{k+1}d(\cdot,x_{0})]$ $\geqslant s_{k}+S(-d(x_{0},\cdot))=s_{k}+\inf_{a\in\mathcal{A}}\sup_{b\in\mathcal{B}}-d(x_{0},T_{ab}(\cdot))\enspace.$
Evaluating the inequality above at point $x_{0}$, we get $s_{k+1}\geqslant s_{k}+\alpha$, which completes the induction. ∎
The case in which there is only one player (Max) is of special interest. Then, the set $\mathcal{A}$ is trivial, being reduced a single action $a$, and we set $T_{b}:=T_{ab}$ for all $b\in\mathcal{B}$. We call such games minimizer-free. The following proposition shows that in this case, the value of the escape rate game coincides with the asymptotic joint displacement introduced by Breuillard and Fujiwara [BF21, Claim 4]:
Proposition 12.
The value of a minimizer-free escape rate game coincides with the following quantity:
(9) $\rho_{\operatorname{disp}}=\lim_{k\to\infty}\inf_{x}\sup_{b_{1},\dots,b_{k}\in\mathcal{B}}\frac{d(T_{b_{k}}\circ\dots\circ T_{b_{1}}(x),x)}{k}\enspace.$
Proof.
In this special case, we have $s_{k}=\sup_{b_{1},\dots,b_{k}\in\mathcal{B}}d(T_{b_{k}}\circ\dots\circ T_{b_{1}}(x_{0}),x_{0})$, and $\rho=\lim_{k}s_{k}/k$. Moreover it is observed in [BF21, Claim 4] that for all choices of $x_{0}$, $\lim_{k}s_{k}/k=\rho_{\operatorname{disp}}$. ∎
3.2. Existence and characterization of the value of the escape rate game
The quantity $\rho$ defined by (7) can be characterized in terms of the non-linear sub-eigenvalues of $S$. We have the following theorem.
Theorem 13.
We have
$\rho=\max\{\lambda\in\mathbb{R}\mid\exists v\in\operatorname{Lip}_{1},\;\lambda+v\leqslant Sv\}\enspace.$
Proof.
We first observe that if $\lambda\in\mathbb{R}$ is such that $\exists v\in\operatorname{Lip}_{1}$, $\lambda+v\leqslant Sv$, then, $k\lambda+v\leqslant S^{k}v$, and since $d(\cdot,x_{0})\geqslant v-v(x_{0})$, $S^{k}d(\cdot,x_{0})\geqslant S^{k}(v-v(x_{0}))=S^{k}v-v(x_{0})$, and so $s_{k}/k\geqslant([S^{k}v](x_{0})-v(x_{0}))/k$, which entails that $\lim_{k}s_{k}/k\geqslant\lambda$. Therefore, $\rho\geqslant\max\{\lambda\in\mathbb{R}\mid\exists v\in\operatorname{Lip}_{1},\;\lambda+v\leqslant Sv\}$.
To prove the reverse inequality, we just need to construct a function $v$ such that $Sv\geqslant\rho+v$. Let $\lambda<\rho$. We first show the following result.
Lemma 14.
For any compact set $C\subset X$, $\exists k_{C}\in\mathbb{N}$ such that $S^{k_{C}}d(\cdot,x_{0})\geqslant k_{C}\lambda+d(\cdot,x_{0})$ on $C$.
Proof.
Let $C$ be a compact subset of $X$. As $d(\cdot,x_{0})\in\operatorname{Lip}_{1}$ and $S$ preserves $\operatorname{Lip}_{1}$, $\forall y\in C$, we have
$[S^{k}d(\cdot,x_{0})](y)$ $\geqslant[S^{k}d(\cdot,x_{0})](x_{0})-d(x_{0},y)$
$=s_{k}-d(x_{0},y)$
$=\lambda k+d(y,x_{0})+(s_{k}-\lambda k)-(d(y,x_{0})+d(x_{0},y))$
$\geqslant\lambda k+d(y,x_{0})+(s_{k}-\lambda k)-\max_{x\in C}(d(x,x_{0})+d(x_{0},x))$
Since $\rho-\lambda>0$ and $\rho=\lim_{k\to\infty}s_{k}/k$, there exists $k_{C}\in\mathbb{N}$ such that $(s_{k_{C}}-\lambda k_{C})-\max_{x\in C}(d(x,x_{0})+d(x_{0},x))\geqslant 0$, which gives the result of the lemma. ∎
We now finish the proof of 13. For convenience, we denote $x\mapsto d(\cdot,x_{0})$ by $w$. Let
(10) $\theta_{C}:=\sup(w,Sw-\lambda,\dots,S^{k_{C}-1}w-(k_{C}-1)\lambda)\enspace.$
As the supremum of nonexpansive functions from $X$ to $\mathbb{R}$, $\theta_{C}$ is also nonexpansive. Moreover, on $C$ we have
(11) $S\theta_{C}$ $\geqslant\sup(Sw,\dots,S^{k_{C}}w-(k_{C}-1)\lambda)$
(12) $\geqslant\sup(Sw,\dots,S^{k_{C}-1}w-(k_{C}-2)\lambda,w+\lambda)=\lambda+\theta_{C}\enspace.$
Let $I$ be the directed set of all compact subsets of $X$ (where the partial order used is the inclusion). We denote $(C_{i})_{i\in I}$, the net containing all compact subsets of $X$, such that $C_{i}=i$. Obviously, $\bigcup_{i\in I}C_{i}=X$ since it contains all singletons $\{x\}$ where $x\in X$. Therefore, the net $(C_{i})_{i\in I}$ converges to $X$. Consider the net $(D_{i})_{i\in I}=(C_{i}\cup\{x_{0}\})_{i\in I}$ of compact sets containing $x_{0}$. Using the same construction as above, we now have a net $(\theta_{D_{i}})_{i\in I}$ verifying $S\theta_{D_{i}}\geqslant\lambda+\theta_{D_{i}}$ on $D_{i}$. We will normalize each $\theta_{D_{i}}$ by subtracting $\theta_{D_{i}}(x_{0})$. The new net $(\theta_{D_{i}})$ is now in $\mathcal{L}_{x_{0}}$, which is compact. So we can extract a converging subnet. We write $\theta_{\lambda}$ its limit. Therefore, as S is continuous, $\forall\lambda<\rho$ we have constructed a map $\theta_{\lambda}\in\mathcal{L}_{x_{0}}$ such that
$S\theta_{\lambda}\geqslant\lambda+\theta_{\lambda}\enspace.$
Taking the limit (up to a subsequence) in the above inequality when $\lambda\to\rho$ we obtain a map $\theta\in\mathcal{L}_{x_{0}}$ verifying $S\theta\geqslant\rho+\theta$. Therefore, $\rho\leqslant\max\{\lambda\in\mathbb{R}\mid\exists v\in\operatorname{Lip}_{1},\;\lambda+v\leqslant Sv\}$. This shows the equality in 13. ∎
Before stating our main theorem, we introduce the notion of $q$-cyclic strategies for player Min. This player chooses a strategy for the game $\Gamma_{k}$ and repeats it indefinitely. We construct it inductively: every $mk+1$ turns, $m\in\mathbb{N}$, player Min forgets the sequence of actions and states played during the $mk$ first turns and then plays the game $\Gamma_{k}$ as if it were starting from $x_{0}$ again. This means that for $\ell=1,\dots,k,m\in\mathbb{N}$ the sequence of actions and states player Min considers to choose their action at turn $mk+\ell$ is
$(x_{0},a_{mk+1},b_{mk+1},T_{a_{mk+1}b_{mk+1}}(x_{0}),$$\dots$
$\!\!\!\!\!\!a_{mk+\ell-1},b_{mk+\ell-1},T_{a_{mk+\ell-1}b_{mk+\ell-1}}\circ\dots\circ T_{a_{mk+1}b_{mk+1}}(x_{0}))$
Note that the infinite sequence of actions played by Min in such a $q$-cyclic strategy will not be periodic unless player Max also selects actions in a cyclic way.
Theorem 15.
The value of the escape rate game exists and is equal to $\rho$. Furthermore, given a sub-eigenvector $v$ associated to $\rho$, player Max possesses a positional stationary optimal strategy obtained by maximizing $b\mapsto v(T_{a_{k+1}b}(x_{k}))$ at turn $k$. Player Min possesses a history dependent optimal strategy, and for all $\epsilon>0$, player Min admits an $\epsilon$-optimal strategy that is $q$-cyclic for some $q$.
Proof.
For the first assertion, we need to show that the two players can guarantee the value $\rho$. For this purpose, we will construct optimal strategies for both players, and in particular a stationary optimal strategy for Max as described in the second assertion.
From 13, there exists $v\in\operatorname{Lip}_{1}$ such that $\rho+v\leqslant Sv$.
Recall that any strategies $\sigma$ and $\tau$ of the two players Min and Max respectively define a sequence $(a_{k},b_{k})_{k\geqslant 0}$ of actions of the two players and that, starting from $x_{0}$, the sequence of states of the game is defined inductively as follows:
$x_{\ell}=T_{a_{\ell}b_{\ell}}(x_{\ell-1})\enspace,\quad\ell\geqslant 1\enspace.$
Let $\sigma$ be an arbitrary strategy of player Min. If $a_{1},b_{1},\dots,b_{k-1}$ is the history of the game at stage $k-1$, the state $x_{k-1}$ of the game is given by
$x_{k-1}\coloneqq T_{a_{k-1}b_{k-1}}\circ\dots\circ T_{a_{1}b_{1}}(x_{0})\enspace,$
and $a_{k}$ is determined by $\sigma$ and the history. We then define the strategy $\tau^{*}$ of player Max, as follows: at step $k$, we select any action
$b_{k}\in\operatorname*{arg\,max}_{b\in\mathcal{B}}v(T_{a_{k}b}(x_{k-1}))\enspace.$
This is a stationary strategy as it depends only on the current state and the last action of Min. We get that the sequence of states generated by the strategies $\sigma$ and $\tau^{*}$ satisfies
$Sv(x_{k-1})=\min_{a\in\mathcal{A}}\max_{b\in\mathcal{B}}v(T_{ab}(x_{k-1}))\leqslant\max_{b\in\mathcal{B}}v(T_{a_{k}b}(x_{k-1}))=v(T_{a_{k}b_{k}}(x_{k-1}))=v(x_{k})\enspace.$
Since $\rho+v\leqslant Sv$, we deduce that $\rho+v(x_{k-1})\leqslant v(x_{k})$, for all $k\geqslant 1$, and by induction
$\rho\leqslant\frac{v(x_{k})-v(x_{0})}{k}\leqslant\frac{d(x_{k},x_{0})}{k}\enspace.$
It follows that the lower value of the game, defined as
$V^{-}\coloneqq\sup_{\tau}\inf_{\sigma}J(\sigma,\tau)$
satisfies
$V^{-}\geqslant\inf_{\sigma}J(\sigma,\tau^{*})\geqslant\rho\enspace.$
So the lower value of the game (and hence the value if it exists) is greater than or equal to $\rho$.
Actually, we can also consider the non zero-sum game where Min still wishes to minimize $J(\sigma,\tau)$ but Max wishes to maximize $J^{-}\coloneqq\liminf_{k\to\infty}J_{k}(a_{1}b_{1}\dots a_{k}b_{k})/k$. In this situation the proof above gives us the stronger inequality
(13) $\sup_{\tau}\inf_{\sigma}\liminf_{k}J_{k}(\sigma,\tau)/k\geqslant\inf_{\sigma}\liminf_{k}J_{k}(\sigma,\tau^{*})/k\geqslant\rho\enspace.$
Now for the upper value, we shall construct for all $\epsilon>0$ a $q$-cyclic $\epsilon$-optimal strategy for Min. Let $\epsilon>0$, as $\rho=\inf_{k\in\mathbb{N}^{*}}\frac{s_{k}}{k}$, there exists $q\in\mathbb{N}^{*}$ such that $\frac{s_{q}}{q}\leqslant\rho+\epsilon$, the value of the game in horizon $q$ exists and is equal to $s_{q}$. Let $\sigma_{q}$ be an optimal strategy of Min for the game $\Gamma_{q}$. We define $\sigma_{q}^{\infty}$ as the $q$-cyclic strategy of Min which consists of applying repeatedly $\sigma_{q}$. Let $\tau\in\mathcal{B}^{s}$. As usual, we denote $(a_{i},b_{i})_{i\in\mathbb{N}^{*}}$ the sequence of actions induced by the strategies $\sigma_{q}^{\infty}$ and $\tau$ and for all $m\in\mathbb{N}$, we define the operator played from turn $(m-1)q+1$ to turn $mq$, $\widetilde{T}_{w_{m}}\coloneqq T_{a_{mq}b_{mq}}\circ\dots\circ T_{a_{(m-1)q+1}b_{(m-1)q+1}}$. As $\sigma_{q}$ is an optimal strategy for the game $\Gamma_{q}$ we have for all $m\in\mathbb{N}$
$d(\widetilde{T}_{w_{m}}(x_{0}),x_{0})\leqslant s_{q}$
So using the properties of $d$ and the nonexpansiveness of the operators we get
$J_{mq}(\sigma_{q}^{\infty},\tau)$ $=d(\widetilde{T}_{w_{m}}\circ\dots\circ\widetilde{T}_{w_{1}}(x_{0}),x_{0})$
$\leqslant d(\widetilde{T}_{w_{m}}\circ\dots\circ\widetilde{T}_{w_{1}}(x_{0}),\widetilde{T}_{w_{m}}(x_{0}))+d(\widetilde{T}_{w_{m}}(x_{0}),x_{0})$
$\leqslant d(\widetilde{T}_{w_{m-1}}\circ\dots\circ\widetilde{T}_{w_{1}}(x_{0}),x_{0})+d(\widetilde{T}_{w_{m}}(x_{0}),x_{0})$
$\vdots$
$\leqslant d(\widetilde{T}_{w_{m}}(x_{0}),x_{0})+\dots+d(\widetilde{T}_{w_{1}}(x_{0}),x_{0})\leqslant ms_{q}\enspace.$
With the same arguments, we have for $0\leqslant\ell<q$,
$J_{mq+\ell}(\sigma_{q}^{\infty},\tau)$ $\leqslant d(T_{a_{mq+\ell}b_{mq+\ell}}\circ\dots\circ T_{a_{mq+1}b_{mq+1}}(x_{0}),x_{0})+ms_{q}$
$\leqslant d(T_{a_{mq+\ell}b_{mq+\ell}}(x_{0}),x_{0})+\dots+d(T_{a_{mq+1}b_{mq+1}}(x_{0}),x_{0})+ms_{q}$
$\leqslant\ell C_{0}+ms_{q}\enspace.$
where $C_{0}$ is an real upper bound of $d(T_{ab}(x_{0}),x_{0})$ for $a\in\mathcal{A}$ and $b\in\mathcal{B}$ which exists by 7. Dividing by $mq+\ell$ and taking the limit superior when $m$ goes to infinity, we can see that we have constructed a q-cyclic strategy for Min such that $\forall\tau$, $J(\sigma_{q}^{\infty},\tau)\leqslant\frac{s_{q}}{q}\leqslant\rho+\epsilon$.
To sum up,
$\forall\epsilon>0,\;\exists(\sigma_{\epsilon}^{*},\tau^{*})\in\mathcal{A}^{s}\times\mathcal{B}^{s}\text{ such that }\rho\leqslant J(\sigma_{\epsilon},\tau^{*})\leqslant\rho+\epsilon\enspace.$
Therefore, the value of the game is $\rho$, Max possesses a positional stationary optimal strategy $\tau^{*}$, and Min possesses cyclic $\epsilon$-optimal strategies.
Now, to construct an optimal strategy for Min, we follow the construction of a cyclic strategy but instead of keeping a fixed period, we increase it by $1$ every time. This means that for the first turn Min plays an optimal strategy associated with $\Gamma_{1}$. Then, he forgets the actions played so far (the other player can remember them!) and plays an optimal strategy for the game $\Gamma_{2}$. Then, he forgets the first turn, and for the $2$ following turns, plays an optimal strategy associated with $\Gamma_{2}$. He repeats this operation, i.e after the $1+\dots+k$ first turns, he forgets the past and plays an optimal strategy associated with $\Gamma_{k+1}$. Denote by $\sigma^{*}$ this strategy of Min. By the same arguments as above, we obtain for every $k\in\mathbb{N}^{*}$, $0\leqslant\ell\leqslant k$, and $\tau\in\mathcal{B}^{s}$ the following property:
$J_{1+\dots+k+\ell}(\sigma^{*},\tau)\leqslant\ell C_{0}+s_{1}+\dots+s_{k}\enspace.$
Dividing by $1+\dots+k+\ell$, we deduce
$\frac{J_{1+\dots+k+\ell}(\sigma^{*},\tau)}{1+\dots+k+\ell}\leqslant\frac{kC_{0}+s_{1}+\dots+s_{k}}{1+\dots+k}\leqslant\frac{2C_{0}}{k+1}+\frac{s_{1}+\dots+s_{k}}{1+\dots+k}\enspace.$
Since $\lim_{k\to\infty}s_{k}/k=\rho$, we deduce that the right hand side converges towards $\rho$. Indeed
$\frac{s_{1}+\dots+s_{k}}{1+\dots+k}\leqslant\frac{s_{1}+\dots+s_{k_{0}}}{1+\dots+k}+\max\{\frac{s_{n}}{n},n\geqslant k_{0}\}\enspace.$
Fix $k_{0}$ and take the limit when $k$ goes to infinity, we obtain
$\limsup_{k\to\infty}\frac{s_{1}+\dots+s_{k}}{1+\dots+k}\leqslant\max\{\frac{s_{n}}{n},n\geqslant k_{0}\}\enspace.$
Now taking the limit when $k_{0}$ tends to infinity, we deduce that
$\limsup_{k\to\infty}\frac{s_{1}+\dots+s_{k}}{1+\dots+k}\leqslant\rho\enspace.$
So
$\limsup_{k\to\infty}\frac{J_{1+\dots+k+\ell}(\sigma^{*},\tau)}{1+\dots+k+\ell}\leqslant\rho\enspace.$
We conclude that
$J(\sigma^{*},\tau)\leqslant\rho$
Remark 2.
Let us consider the minimizer-free matrix multiplication game, and suppose that the matrices are invertible. Then, we can find $v\in\operatorname{Lip}_{1}$ verifying $Sv\leqslant\rho+v$ such that the function $V=e^{v}$ is convex (and positively homogeneous of degree $1$) (this is straightforward refinement of the proof of 13, noting that in the minimizer-free case, the operator $S$ preserves the set of functions $v$ such that $V=e^{v}$ is convex).Then, by considering $W(x)\coloneqq\sqrt{V(xx^{*})}$ we get that $W$ is a norm verifying $S(W^{2})\geqslant\exp(\rho)W^{2}$. This is reminiscent of the characterization of the joint spectral radius in terms of extremal norms (see [JM16] – note however that the order is reversed here).
A direct corollary of this theorem is the existence of a uniform value as defined by Mertens and Neyman [MN81].
Definition 16 (Uniform Value).
A zero-sum repeated game is said to have a uniform value $v_{\infty}$ if both players can guarantee $v_{\infty}$ up to an arbitrary precision provided that they play for a sufficiently long time. Formally $v_{\infty}$ is the uniform value of the game if $\forall\epsilon>0$, there exist a couple of strategies $(\sigma_{\epsilon},\tau_{\epsilon})$ and a time $N$ such that $\forall n\geqslant N$ and $\forall(\sigma,\tau)$
$\frac{J_{n}(\sigma_{\epsilon},\tau)}{n}\leqslant v_{\infty}+\epsilon$
$\frac{J_{n}(\sigma,\tau_{\epsilon})}{n}\geqslant v_{\infty}-\epsilon$
See [MN81] for more information on the uniform value, see also [BGV14].
Corollary 17.
$\rho$ is the uniform value of the game $\Gamma$.
Proof.
The strategy $\tau^{*}$ defined in the proof of 15 gives us the second inequality for all $\epsilon$ and for all $N$. Because there exist an optimal strategy $\sigma^{*}$ for Min, i.e $\forall\tau\in\mathcal{B}^{s},\limsup_{k\to\infty}\frac{J_{k}(\sigma^{*},\tau)}{k}\leqslant\rho$, the first inequality is obviously verified if $N\in\mathbb{N}^{*}$ is big enough. ∎
Then we have the direct corollaries for vector addition games and matrix multiplication games.
Corollary 18.
Vector addition games have a uniform value
Corollary 19.
Monotone and Invertible matrix multiplication games have a uniform value.
For strictly positive matrix multiplication games, we can also obtain an optimal strategy for Min, using the ideas around tubular sets from [BM01]. A proof can be found in section 3.4.
3.3. Distance-like functions
Recall that we say a $1$-Lipschitz function $v:X\to\mathbb{R}$ is distance-like if, there exist $x_{0}\in X$ and a constant $\alpha\in\mathbb{R}$ such that
$v(x)\geqslant\alpha+d(x,x_{0}),\quad\forall x\in X\enspace.$
Distance-like functions allow one to bound the value from above.
Proposition 20.
Suppose that there exists a distance-like function $v$ such that
$Sv\leqslant\lambda+v\enspace,$
for some $\lambda\in\mathbb{R}$. Then, the value $\rho$ of the escape rate game satisfies $\rho\leqslant\lambda$. Moreover, by selecting at state $x$ an action $a$ which achieves the minimum in the expression of $Sv(x)$, we obtain a stationary strategy for player Min which guarantees $\lambda$.
Proof.
By using this strategy, we get (by a immediate induction) that for any sequence of actions $(b_{k})_{n\in\mathbb{N}^{*}}$ played by Max, setting $x_{k}=T_{a_{k}b_{k}}\circ\dots\circ T_{a_{1}b_{1}}(x_{0})$,
$d(x_{k},x_{0})+\alpha$ $\leqslant v(x_{k})=v(T_{a_{k}b_{k}}x_{k-1})$
$\leqslant\sup_{b\in\mathcal{B}}v(T_{a_{k}b}x_{k-1})=Sv(x_{k-1})$
$\leqslant v(x_{k-1})+\lambda\leqslant\cdots$
$\leqslant v(x_{0})+k\lambda\enspace,$
showing that player Min can guarantee $\lambda$. Hence, the value of the escape rate game satisfies $\rho\leqslant\lambda$. ∎
We shall say that a nonexpansive map $T:X\to X$ is an almost-isometry if for some $\gamma\geqslant 0$,
(14) $d(T(x),T(y))\geqslant d(x,y)-\gamma\enspace,\forall x,y\in X\enspace.$
In particular, an isometry is an almost-isometry (with $\gamma=0$). We shall say that the nonexpansive maps $T_{ab}$ with $(a,b)\in A\times B$ constitute a uniform family of almost-isometries if every map $T=T_{ab}$ satisfies (14) for some constant $\gamma\geqslant 0$ independent of the choice of $(a,b)\in A\times B$.
Proposition 21.
Suppose that the maps $T_{ab}$ with $(a,b)\in A\times B$ constitute a uniform family of almost-isometries. Then, the Shapley operator $S$ preserve the space of distance-like functions $\mathscr{D}$.
Proof.
Since $S$ commutes with the addition of a constant, it suffices to show that $Sd(\cdot,x_{0})$ is distance like. Let $M:=\max_{(a,b)\in A\times B}d(x_{0},T_{ab}(x_{0}))$, and observe that $M<\infty$ thanks to 7. Then, for all $(a,b)\in A\times B$,
$d(T_{ab}(x),x_{0})\geqslant d(T_{ab}(x),T_{ab}(x_{0}))-d(x_{0},T_{ab}(x_{0}))\geqslant d(x,x_{0})-\gamma-M,\qquad\forall x\in\mathbb{R}^{n}\enspace.$
This implies that $Sd(\cdot,x_{0})\geqslant d(\cdot,x_{0})-\gamma-M$. ∎
We now prove the dual characterization of the value of the game, stated as 5 in the introduction.
Proof of 5.
It follows from 20 that $\rho\leqslant\inf\{\lambda\in\mathbb{R}\mid v\in\mathscr{D},Sv\leqslant\lambda+v\}$. The proof of the reverse inequality relies on the following lemma.
Lemma 22.
For all $\lambda>\rho$ there exists $k\in\mathbb{N}^{*}$ such that $S^{k}d(\cdot,x_{0})\leqslant k\lambda+d(\cdot,x_{0})$.
Proof.
We have $S^{k}d(\cdot,x_{0})\in\operatorname{Lip}_{1}$ so
$S^{k}d(\cdot,x_{0})\leqslant[S^{k}d(\cdot,x_{0})](x_{0})+d(\cdot,x_{0})=s_{k}+d(\cdot,x_{0})\enspace.$
Since $\inf_{k\geqslant 1}s_{k}/k=\rho$ and $\lambda-\rho>0$, there exists $k\geqslant 1$ such that $s_{k}/k\leqslant\lambda$. Then, $Sd(\cdot,x_{0})\leqslant k\lambda+d(\cdot,x_{0})$. ∎
The end of the proof relies on a dual version of the “sup-averaging” argument (10) used in the proof of 13. We take $k$ as in (22), and define
$\theta\coloneqq\inf\{d(\cdot,x_{0}),Sd(\cdot,x_{0})-\lambda,\dots,S^{k-1}d(\cdot,x_{0})-(k-1)\lambda\}\enspace.$
Since, by 21, $S$ preserves the set of distance-like functions, $\theta$ is distance-like. Moreover we have
$S\theta$ $\leqslant\inf\{Sd(\cdot,x_{0}),S^{2}d(\cdot,x_{0})-\lambda,\dots,S^{k}d(\cdot,x_{0})-(k-1)\lambda\}$
$\leqslant\inf\{Sd(\cdot,x_{0}),S^{2}d(\cdot,x_{0})-\lambda,\dots,S^{k-1}d(\cdot,x_{0})-(k-2)\lambda,d(\cdot,x_{0})+\lambda\}$
$=\theta+\lambda\qed$
The infimum in 5 is not reached in all generality, i.e for some families of isometries there cannot exist a distance-like function such that $Sv\leqslant\rho+v$ as proven by the following example.
Example 23.
Take $X$ as being the Poincaré half space $\{z\in\mathbb{C}\mid\Im(z)>0\}$ equipped with the Poincaré Metric
$d(z,z^{\prime})=2\operatorname{arcsinh}(\frac{\lVert z-z^{\prime}\rVert}{\sqrt{\Im(z)}\sqrt{\Im(z^{\prime})}})\enspace,$
where $\Im(\cdot)$ denotes the imaginary part of a complex number. We suppose the two players are dummies, so that a single dynamics is available $T(z)=z+1$. It is trivial to see that $\lim_{k\to\infty}\frac{d(T^{k}(i),i)}{k}=0$. Then if there exists a distance-like function $v$ (of constant $\gamma\in\mathbb{R}$ and center $x_{0}$) such that for all $x\in X$, we have $v(Tx)\leqslant v(x)$. Then we would have for all $k\in\mathbb{N}^{*}$
$\gamma+2\frac{\operatorname{arcsinh}(k)}{\Im(x_{0})}\leqslant v(T^{k}(x_{0}))\leqslant v(x_{0})$
which is absurd as $\lim_{k\to\infty}\operatorname{arcsinh}(k)=+\infty$.
Combining 20 and 5, we obtain:
Corollary 24.
Suppose that the maps $T_{ab}$ with $(a,b)\in A\times B$ constitute a uniform family of almost-isometries. Then, for all $\lambda>\rho$, player Min has a stationary strategy which guarantees $\lambda$. ∎
3.4. Dual characterization of the value for cones equipped with the Funk metric
Consider a closed pointed convex cone $C\subset\mathbb{R}^{n}$. We define a cross section of $C$ by $\Delta\coloneqq\{x\in C\mid\langle x,e^{*}\rangle=1\}$ where $e^{*}$ is a point in the interior of the dual cone $C^{*}$.
Lemma 25.
$\Delta$ is compact for the usual Euclidean topology
Proof.
Indeed, as the preimage of $\{1\}$ by $e^{*}$, it is closed. Moreover, if it were not bounded, this would mean that there exists $(x_{k})_{k\in\mathbb{N}}\in\Delta^{\mathbb{N}}$ such that $\lVert x_{k}\rVert\to\infty$. The sequence $y_{k}\coloneqq\frac{x_{k}}{\lVert x_{k}\rVert}$ lies on the unit sphere, which is compact, so it converges, up to a subsequence, to a certain $y$ in the intersection of the unit sphere and $C$ because it is closed. However $(x_{k})_{k\in\mathbb{N}}\in\Delta^{\mathbb{N}}$ so $\langle y,e^{*}\rangle=0$ so $y=0$ which is absurd as it is on the unit sphere. So $\Delta$ is bounded and therefore compact. ∎
[…]
We introduced $\Delta$ as for all $\lambda\in\mathbb{R}_{>0},(x,y)\in\operatorname{int}C,\operatorname{Funk}(\frac{x}{\lambda},y)=\operatorname{Funk}(x,y)-\log(\lambda)$. So the core idea in this section is that, up to a renormalization by $\log(\langle T_{a_{k}b_{k}}\circ\dots T_{a_{1}b_{1}}(x_{0}),e^{*}\rangle)$, the game played on $relint\Delta$ is the same. So, if we assumed that the operators $T_{ab}$ can be continuously extended (considering the Euclidean topology) on $C$, it seems natural to consider a traditional game played on $\Delta$ with the normalized operators $\hat{T}_{ab}(x)=\frac{T_{ab}(x)}{\langle T_{ab}(x),e^{*}\rangle}$ and paiement at turn $k$ equal to $\log(\langle T_{a_{k}b_{k}}(\hat{x}_{k-1}),e^{*}\rangle)$. Where $\hat{x_{k}}\coloneqq\hat{T}_{a_{k}b_{k}}\circ\dots\circ\hat{T}_{a_{1}b_{1}}(x_{0})$. Indeed the payoff of the game in horizon $k$ would be
$\sum_{k=1}^{n}\log(\langle T_{a_{k}b_{k}}(\hat{x}_{k-1}),e^{*}\rangle)=\log(\langle T_{a_{n}b_{n}}\circ\dots\circ T_{a_{1}b_{1}}(x_{0}),e^{*}\rangle)$
We now introduce the Shapley operator of this game, operating on $\mathscr{C}(\Delta)$
$Fv(x)\coloneqq\inf_{a\in\mathcal{A}}\sup_{b\in\mathcal{B}}\Big{\{}\log(\langle T_{ab}(x),e^{*}\rangle)+v\big{(}\frac{T_{ab}(x)}{\langle T_{ab}(x),e^{*}\rangle}\big{)}\Big{\}}$
Lemma 26.
Assume that the map $(a,b,x)\mapsto T_{ab}(x)$ is continuous, then F preserves $\mathscr{C}(\Delta)$.
Proof.
This comes from the compactness of $\Delta,\mathcal{A}$ and $\mathcal{B}$. Since linear maps between finite dimensional space are continuous
$f_{a}:x\mapsto\sup_{b\in\mathcal{B}}\phi(x,b)\coloneqq\sup_{b\in\mathcal{B}}\Big{\{}\log(\langle T_{ab}(x),e^{*}\rangle)+v\big{(}\frac{T_{ab}(x)}{\langle T_{ab}(x),e^{*}\rangle}\big{)}\Big{\}}$
is lower semi-continuous as the supremum of continuous function. Let $x^{*}\in\Delta$ and $(x_{n})_{n\in\mathbb{N}}$ be a sequence converging to $x^{*}$. Because $\mathcal{B}$ is compact, $\exists b^{*}\in\mathcal{B}$ such that $f_{a}(x^{*})=\phi(x^{*},b^{*})$.
Then
$\limsup_{n\to\infty}f_{a}(x_{n})\leqslant\limsup_{n\to\infty}\phi(x_{n},a^{*})=f_{a}(x^{*})$
So $f_{a}$ is continuous. The rest of the proof is similar, using the joint continuity of $(a,b,x)\mapsto T_{ab}(x)$. ∎
So the space $\mathscr{C}\coloneqq\mathscr{C}(\Delta)\cap\operatorname{Lip}_{1}(\Delta)$ is also preserved by $F$.
This space equipped with the hemi-metric $\delta:(f,g)\mapsto\max_{x\in\Delta}f(x)-g(x)$ is a complete metrically star shaped hemi-metric space. For completeness we recall the definition of a metrically star-shaped space, from [Vig13].
Definition 27.
We say that $(X,d)$ is metrically star-shaped with center $x^{*}$ if there exists a family of geodesics $(\gamma_{y})_{y\in X}$, such that $\gamma_{y}$ joins $x^{*}$ to $y$ and such that the following inequality is satisfied for every $(y,z)\in X^{2}$ and $s\in[0,1]$:
$d(\gamma_{y}(s),\gamma_{z}(s))\leqslant sd(y,z)$
If $d$ is only a hemi-metric, we additionally require that for any $y$, $\lim_{s\to 1}d(y,\gamma_{y}(s))=0$
$F$ is nonexpansive with respect to $\delta$, therefore theorem 3 tells us that
$\bar{\rho}\coloneqq\lim_{k\to\infty}\frac{\delta(T^{k}(0),0)}{k}=\lim_{k\to\infty}\frac{\max_{x\in\Delta}F^{k}(0)(x)}{k}$ $=\inf_{v\in\mathcal{C}}\max_{x\in\Delta}F(v)(x)-v(x)$
$=\inf\{\lambda\mid v\in\mathscr{C},Fv\leqslant\lambda+v\}$
Now, we will see that $\bar{\rho}=\rho$. For that, we need the following lemma.
Lemma 28.
There exists $K\in\mathbb{R}^{+}$ such that for all $x\in C/\{0\}$
$\operatorname{Funk}(x,x_{0})-K\leqslant\log(\langle x,e^{*}\rangle)\leqslant\operatorname{Funk}(x,x_{0})+K$
Proof.
By the Hahn Banach separation theorem, $x\leqslant y$, $(x,y)\in C^{2}$ if and only if $\langle x,u\rangle\leqslant\langle y,u\rangle$ for all $u\in C^{*}$. So for all $(x,y)\in C\times\operatorname{int}(C)$
$\operatorname{Funk}(x,y)=\log\sup_{u\in C^{*}}\frac{\langle x,u\rangle}{\langle y,u\rangle}$
So $\operatorname{Funk}(x,x_{0})\geqslant\log\frac{\langle x,e^{*}\rangle}{\langle x_{0},e^{*}\rangle}=\log(\langle x,e^{*}\rangle)-\log(\langle x_{0},e^{*}\rangle)$ For the other inequality, the map $x\mapsto\operatorname{Funk}(x,x_{0})$ is continuous for the Euclidean topology so the inequality is verified for $x\in\Delta$. If it verified in $\Delta$ it is also true for $x\in C$. Indeed
$\operatorname{Funk}(x,x_{0})$ $=\operatorname{Funk}(\frac{x}{\langle x,e^{*}\rangle},x_{0})+\log(\langle x,e^{*}\rangle)$
$\leqslant\log(\langle\frac{x}{\langle x,e^{*}\rangle},e^{*}\rangle)+K+\log(\langle x,e^{*}\rangle)=\log(\langle x,e^{*}\rangle)+K$
This "equivalence" between $Funk(x,x_{0})$ and and $\log(\langle x,e^{*}\rangle)$ tells us that both games have the same value as long as the starting point is in the interior of $C$. So, in particular $\lim_{k\to\infty}\frac{F^{k}(0)(x_{0})}{k}\leqslant\bar{\rho}$, so $\rho\leqslant\bar{\rho}$. For the other inequality we use corollary 23 from [GV12]. It tells us that there exists $x^{*}\in\Delta$ such that for all $k\in\mathbb{N}$
$\frac{F^{k}(0)(x^{*})}{k}\geqslant\bar{\rho}$
Therefore we can construct a strategy for the maximizer to guarantee $\bar{\rho}$ if the starting point is $x^{*}$. If $x^{*}$ lies on the border of the cone the lower bound of the following lemma is not usefulm however, we still have for all sequence of actions $(a_{n},b_{n})_{n\in\mathbb{N}}$ and all $k\in\mathbb{N}^{*}$ that
$\log(\langle T_{a_{k}b_{k}}\circ\dots T_{a_{1}b_{1}}(x^{*}),e^{*}\rangle)\leqslant\operatorname{Funk}(T_{a_{k}b_{k}}\circ\dots T_{a_{1}b_{1}}(x^{*}),x_{0})+K$
$\leqslant\operatorname{Funk}(T_{a_{k}b_{k}}\circ\dots T_{a_{1}b_{1}}(x^{*}),T_{a_{k}b_{k}}\circ\dots T_{a_{1}b_{1}}(x_{0}))+\operatorname{Funk}(T_{a_{k}b_{k}}\circ\dots T_{a_{1}b_{1}}(x_{0}),x_{0})+K$
$\leqslant\operatorname{Funk}(x^{*},x_{0})+\operatorname{Funk}(T_{a_{k}b_{k}}\circ\dots T_{a_{1}b_{1}}(x_{0}),x_{0})+K$
As $\operatorname{Funk}(x^{*},x_{0})$ is finite, an optimal strategy for Min in the original game will guarantee $\rho$ in the game played on $\Delta$. Therefore $\rho\geqslant\bar{\rho}$ and the equality is verified. Finally, we can identify $\mathscr{C}$ to the set of continuous (for the Euclidean topology) and distance-like function on $C$, by posing $v(\lambda x)=\log(\lambda)+v(x)$ for all $(\lambda,x)\in\mathbb{R}\times\Delta$. We $\mathscr{C}_{X}$ this set. For such functions it is equivalent to verify $Sv\leqslant\lambda+v$ on $C$ and $Fv\leqslant\lambda+v$ on $\Delta$. So we get the following dual characterization
Theorem 29.
The value $\rho$ of escape rate games played on cones has the following dual characterization
$\rho=\max\{\lambda\in\mathbb{R}\mid\exists v\in\operatorname{Lip}_{1},Sv\geqslant\lambda+v\}=\inf\{\lambda\in\mathbb{R}\mid\exists v\in\mathscr{C}_{X},Sv\leqslant\lambda+v\}$
Unfortunately, there is no guarantee that there exists a function $v\in\mathscr{C}_{X}$ verifying $Sv\leqslant\rho+v$ (or $Fv\leqslant\rho+v$). However, the above section tells us that if the "2-players cohomological equation"
$Fv=\lambda+v$
has a solution $(\mu,v)\in\mathbb{R}\times\mathscr{C}$, then $\mu$ is necessarily equal to $\rho$. Indeed, in such a case the lifted function verifies $Sv=\mu+v$ and we can construct stationary optimal strategies guaranteeing $\mu$ in the game where the payoff is
$\limsup_{k\to\infty}\frac{v(T_{a_{k}b_{k}}\circ\dots\circ T_{a_{1}b_{1}}(x_{0}))}{k}$
by playing $a_{k}\in\operatorname*{arg\,min}_{a\in\mathcal{A}}\max_{b\in\mathcal{B}}v(T_{ab}(x_{k-1}))$ and $b_{k}\in\operatorname*{arg\,max}_{b\in\mathcal{B}}v(T_{a_{k}}b(x_{k-1}))$. Then as $v$ is continuous, it is bounded on $\Delta$ and because $v(\lambda x)=\lambda+v(x)$, for all $x\in\Delta$, we get that there exists $M\in\mathbb{R}^{+}$ such that for all $x\in X$
(15) $\log(\langle x,e^{*}\rangle)-M\leqslant v(x)\leqslant\log(\langle x,e^{*}\rangle)+M$
This inequalities tell us that the two games are equivalent.
Remark 3.
In the traditional cohomological equation one looks to solve the equation
$g=f\circ T-f$
where $g$ is a given function and $(X,T)$ the dynamical system studied.
The existence of such an eigenfunction is guaranteed in the tubular case. As we deal with the Funk Metric, we deal with cones with compact section instead of tubes but we keep the former terminology as the arguments for the following proposition were essentially given by Bousch and Mairesse when considering the 1-player case for topical functions[BM01]. Here, the compactness has to be understand with respect to the Thompson metric because as proved beforehand, the cross-sections considered are always compact with respect to the Euclidean topology. The issue lies within the boundary of $\Delta$ as the Thompson and Euclidean topologies are equivalent on the relative interior of a cross section.
Proposition 30.
If the operators $(T_{ab})$ preserve a cone with a compact section $C$ then there exist a function $v\in\mathscr{C}_{C}$ such that
$Fv=\rho+v$
Proof.
We saw in 10 that an eigenvalue $\lambda\in R$ exists when considering the $\operatorname{Lip}_{1}$ functions and lipschitz function for the Funk metric are additively log-homogeneous. So, if the operators $(T_{ab})$ preserve a cone with compact section $C$, we only need to consider the game on $C$ and proposition 10 ensures the existence of a continuous and log-homogeneous function verifying inequality 15 as the cross section is compact. Then, $\lambda$ is the value of the game so $\lambda=\rho$. ∎
4. Application to Vector Addition Games
As an illustration, we give a complete solution of the vector addition game. We denote by $((\mathbb{R}^{n})^{*},\|\cdot\|_{*})$ the dual normed space of $(\mathbb{R}^{n},\|\cdot\|)$. Let us first notice that this game doesn’t not depend on the initial state and is trivial if $-\mathcal{B}\subset\mathcal{A}$. Indeed for every $k\in\mathbb{N}^{*}$, Min would just need to play $-b_{k}$ at turn $k+1$ and then the game would essentially stay in the state $x_{0}+a_{1}$ and therefore the escape rate would be equal to zero.
Proposition 31.
The game has a value equal to
(16) $\lambda\coloneqq\max_{b\in\mathcal{B}}\min_{a\in\operatorname*{co}(\mathcal{A})}\lVert a+b\rVert$
Any linear form $\ell$ achieving the maximum in
$\max_{\ell\in(\mathbb{R}^{n})^{*},\|\ell\|_{*}\leqslant 1}\max_{b\in\mathcal{B}}\min_{a\in\operatorname*{co}(\mathcal{A})}\ell(a+b)$
is a solution of the eigenproblem $S\ell=\lambda+\ell$. Furthermore, Max has a constant optimal strategy by playing
(17) $b^{*}\in\operatorname*{arg\,max}_{b\in\mathcal{B}}\min_{a\in\operatorname*{co}(\mathcal{A})}\lVert a+b\rVert$
Finally, the function
$\phi(x)=\text{dist}(x,-(n+1)\operatorname*{co}(\mathcal{A}))\enspace.$
where $-n\operatorname*{co}(\mathcal{A})$ has to be understood as a Minkowski sum, satisfies $S\phi\leqslant\lambda+\phi$. By playing at state $x\in\mathbb{R}^{n}$ an action $a$ which achieves the minimum in the expression of $S\phi(x)$, we obtain a stationary optimal strategy of Min.
Proof.
We have
$\lambda$ $=\max_{b\in\mathcal{B}}\min_{a\in\operatorname*{co}(\mathcal{A})}\lVert a+b\rVert=\max_{b\in\mathcal{B}}\min_{a\in\operatorname*{co}(\mathcal{A})}\max_{\ell\in(\mathbb{R}^{n})^{*},\|\ell\|_{*}\leqslant 1}\ell(a+b)$
$=\max_{b\in\mathcal{B}}\max_{\ell\in(\mathbb{R}^{n})^{*},\|\ell\|_{*}\leqslant 1}\min_{a\in\operatorname*{co}(\mathcal{A})}\ell(a+b)\ =\max_{\ell\in(\mathbb{R}^{n})^{*},\|\ell\|_{*}\leqslant 1}\max_{b\in\mathcal{B}}\min_{a\in\operatorname*{co}(\mathcal{A})}\ell(a+b)$
Let $\ell$ be a linear form achieving the maximum in the expression above. Then we get
$S\ell$ $=\ell+\min_{a\in\mathcal{A}}\max_{b\in\mathcal{B}}\ell(a+b)=\ell+\min_{a\in\operatorname*{co}(\mathcal{A})}\max_{b\in\operatorname*{co}(\mathcal{B})}\ell(a+b)$
$=\ell+\max_{b\in\operatorname*{co}(\mathcal{B})}\min_{a\in\operatorname*{co}(\mathcal{A})}\ell(a+b)=\ell+\max_{b\in\mathcal{B}}\min_{a\in\operatorname*{co}(\mathcal{A})}\ell(a+b)$
$=\ell+\max_{\ell\in(\mathbb{R}^{n})^{*},\|\ell\|_{*}\leqslant 1}\max_{b\in\mathcal{B}}\min_{a\in\operatorname*{co}(\mathcal{A})}\ell(a+b)=\ell+\lambda$
Let Max play $b^{*}\in\operatorname*{arg\,max}_{b\in\mathcal{B}}\min_{a\in\operatorname*{co}(\mathcal{A})}$ then for any $k\in\mathbb{N}^{*}$ we have $\forall(a_{1},\dots,a_{k})\in\mathcal{A}^{k}$
$\frac{\lVert x_{0}+a_{1}+b^{*}+\dots+a_{k}+b^{*}-x_{0}\rVert}{k}=\lVert\frac{a_{1}+\dots+a_{k}}{k}+b^{*}\rVert\geqslant\max_{b\in\mathcal{B}}\min_{a\in\operatorname*{co}(\mathcal{A})}\lVert a+b\rVert$
since $\frac{a_{1}+\dots a_{k}}{k}\in\operatorname*{co}(\mathcal{A})$
The above equalities are obtain using Sion minimax theorem ([Sio58]) and the characterization of the norm as a supremum over linear forms.
For the second inequality, let us first recall the useful Shapley Folkman lemma.
Lemma 32 (Shapley Folkman [Sta69]).
Let $N\in\mathbb{N}^{*}$ such that $N>n$. If $X_{1},\dots,X_{N}$ are nonempty bounded subset of $\mathbb{R}^{n}$ then for any $x\in\operatorname*{co}(\sum_{k=1}^{N}X_{k})=\sum_{k=1}^{N}\operatorname*{co}(X_{k})$, there exists $I\subset\{1,\dots,N\}$ such that $|I|\leqslant n$ and
$x=\sum_{i\in I}q_{i}+\sum_{k\in\{1,\dots,N\}/I}q_{k}$
where $q_{i}\in\operatorname*{co}(X_{i})/X_{i}$ and $q_{k}\in X_{k}$
Let $\tau\in\mathcal{B}^{s}$ a strategy of Max. To simplify the notations, we are going to denote $\max_{b\in\mathcal{B}}\min_{a\in\operatorname*{co}(\mathcal{A})}\lVert a+b\rVert=\max_{b\in\mathcal{B}}d(b,-\operatorname*{co}(\mathcal{A}))$ by $\lambda$. Given a strategy $\sigma$ of Min we will denote the resulting position after $k$ turns, i.e $x_{0}+a_{1}+b_{1}+\dots+a_{k}+b_{k}$ by $x_{k}(\sigma,\tau)$. We are going to prove that if $x_{0}\in-n\operatorname*{co}(\mathcal{A})$, there exists a strategy $\sigma^{*}\in\mathcal{A}^{s}$ such that for all $k\in\mathbb{N}$
$\delta_{k}\coloneqq d(x_{k}(\sigma^{*},\tau),-n\operatorname*{co}(\mathcal{A}))\leqslant k\lambda$
where $n\operatorname*{co}(\mathcal{A})$ as to be understood as a Minkowski sum.
This would mean that, starting from $-n\operatorname*{co}(\mathcal{A})$ we can guarantee a payoff lower than lambda and, as the initial position does not matter, we would have proven our proposition. Given the preceding sentence, we can consider the game where Max plays first.
Let us assume that for a certain $k\in\mathbb{N}$, the above inequality is true. It means that $x_{k}\in-n\operatorname*{co}(\mathcal{A})+k\overline{B(\lambda)}$ and since, by definition, the move $b_{k+1}$ played by Max given the strategy $\tau$ is in $-\operatorname*{co}(\mathcal{A})+\overline{B(\lambda)}$. So
$x_{k}+b_{k+1}\in-(n+1)\operatorname*{co}(\mathcal{A})+(k+1)\overline{B(\lambda)}$
The Shapley-Folkman lemma tells us that $-(n+1)\operatorname*{co}(\mathcal{A})\subset-n\operatorname*{co}(\mathcal{A})-\mathcal{A}$. So there exists at least one $\hat{a}\in\mathcal{A}$ such that
$x_{k}+b_{k+1}\in-n\operatorname*{co}(\mathcal{A})-\hat{a}+(k+1)\overline{B(\lambda)}$
So by playing $a_{k}+1=\hat{a}$, we get that
$x_{k+1}\in-n\operatorname*{co}(\mathcal{A})+(k+1)\overline{B(\lambda)}$
Then our proposition follows by induction. Now, this settings not only helps to better understand what escape rate games are, but it provides a great example for the usefulness of our results as 20 and the fact that $\phi$ is distance-like function gives us the reverse inequality immediately. Indeed
$S\phi(x)$ $=\min_{a\in\mathcal{A}}\max_{b\in\mathcal{B}}d(x+a+b,-(n+1)\operatorname*{co}(\mathcal{A}))=\min_{a\in\mathcal{A}}\max_{b\in\mathcal{B}}\min_{c\in(n+1)\operatorname*{co}(\mathcal{A})}\lVert x+a+b+c\rVert$
$\leqslant\min_{a\in\mathcal{A}}\min_{a^{\prime}\in n\operatorname*{co}(\mathcal{A})}\max_{b\in\mathcal{B}}\min_{d\in\operatorname*{co}(\mathcal{A})}\lVert x+a+a^{\prime}+b+d\rVert$
$\leqslant\min_{a\in\mathcal{A}}\min_{a^{\prime}\in n\operatorname*{co}(\mathcal{A})}\max_{b\in\mathcal{B}}\min_{d\in\operatorname*{co}(\mathcal{A})}\lVert x+a+a^{\prime}\rVert+\lVert b+d\rVert$
$\leqslant\min_{a\in(n+1)\operatorname*{co}(\mathcal{A})}\max_{b\in\mathcal{B}}\min_{d\in\operatorname*{co}(\mathcal{A})}\lVert x+c+a\rVert+\lVert b+d\rVert$
$=\phi(x)+\max_{b\in\mathcal{B}}\min_{d\in\operatorname*{co}(\mathcal{A})}\lVert b+d\rVert$
We get from the third to the fourth line using the Shapley-Folkman lemma. ∎