### v2

# The competitive spectral radius of families of nonexpansive mappings

Abstract.
We consider a new class of repeated zero-sum games in which the payoff is the escape rate of a switched dynamical system, where at every stage, the transition is given by a nonexpansive operator depending on the actions of both players. This generalizes to the two-player (and non-linear) case the notion of joint spectral radius of a family of matrices. We show that the value of this game does exist, and we characterize it in terms of an infinite dimensional non-linear eigenproblem. This provides a two-player analogue of Mañe’s lemma from ergodic control and extends to the two-player case results of Kohlberg and Neyman (1981), Karlsson (2001), and Vigeral and the second author (2012), concerning the asymptotic behavior of nonexpansive mappings. We discuss two special cases of this game: order preserving and positively homogeneous self-maps of a cone equipped with Funk’s and Thompson’s metrics, and translations of a finite dimensional normed space.
Key words and phrases:
Nonexpansive mappings, joint spectral radius, Hilbert and Thompson metrics, zero-sum games, uniform value
1. Introduction
1.1. Context
A self map $T$ of a metric space $(X,d)$ is nonexpansive if $d(Tx,Ty)\leqslant d(x,y)$ for all $x,y\in X$. Then, its escape rate is defined by
$\rho(T)=\lim_{k\to\infty}\frac{d(T^{k}\bar{x},\bar{x})}{k}\enspace,$
where $\bar{x}$ is an arbitrary element of $X$. Since $T$ is nonexpansive, the existence of the limit follows from a subadditive argument, and the limit is independent of the choice of $\bar{x}$. Kohlberg and Neyman proved the following characterization of $\rho(T)$ when $T$ acts on a subset of a normed space.
Theorem 1 ([KoNe]).
Let $X$ be a star-shaped subset of a normed space $E$ and $T:X\to X$ be a nonexpansive map, then there exists a continuous linear form $f$ on $E$ of dual norm one such that, for all $x\in X$,
$\rho(T)=\lim_{k\to\infty}\frac{\lVert T^{k}x\rVert}{k}=\inf_{y\in X}\lVert Ty-y\rVert=\lim_{k\to\infty}\langle f,\frac{T^{k}x}{k}\rangle\enspace.$
Moreover, $f$ can be chosen in the weak-* closure of the set of extreme points of the dual unit ball.
They also showed that for all $x\in X$ it is possible to choose $f$ in 1 such that for all $k\in\mathbb{N}^{*}$, we have
$\langle f,T^{k}x\rangle-\langle f,x\rangle\geqslant k\rho(T)\enspace.$
Karlsson extended this result by proving the following Denjoy-Wolff type theorem, which relies on Gromov’s compactification of a metric space by horofunctions [Gromov1981]. The latter are particular Lipschitz functions of Lipschitz constant $1$ which arise when compactifying the space $X$ as follows: we identify a point $x\in X$ to the function $i(x)\coloneqq d(x,\bar{x})-d(x,\cdot)$ from $X$ to $\mathbb{R}$, where $\bar{x}\in X$ is an arbitrary basepoint; then, we define the horocompactification of $X$ as the closure of $i(X)$ in the topology of pointwise convergence. (This choice of topology was made e.g. in [rieffel, GAUBERT_2011, Karlsson2024]; the topology of uniform convergence on bounded sets was originally considered [Gromov1981], leading to a different class of horofunctions). The functions of the closure $\overline{i(X)}$ are called metric functionals [Karlsson2024]. The functions of $\overline{i(X)}/i(X)$ are called horofunctions, they constitute the horoboundary.
Theorem 2 (See [Ka]).
Let $X$ be a proper metric space, $T$ a nonexpansive self-map of $X$, and suppose that every orbit of $T$ in $X$ is unbounded. Then, for all $x\in X$, there is a horofunction $h$ such that
$h(T^{k}x)-h(x)\geqslant k\rho(T)\quad\forall k\geqslant 0\enspace.$
Gaubert and Vigeral refined 2, showing that the horofunction $h$ can be chosen to be independent of $x\in X$ when $X$ satisfies an assumption of non-positive curvature in the sense of Busemann. They consider more generally maps $T$ which are nonexpansive with regard to a hemi-metric of $X$, i.e. a map $d$ taking values in $\mathbb{R}$, verifying the triangular inequality and a weak separation axiom. They established the following maximin characterization of the escape rate.
Theorem 3 (See [GAUBERT_2011]).
Let $T$ be a nonexpansive self-map of a complete metrically star-shaped hemi metric space $(X,d)$. Then
$\rho(T)=\inf_{y\in X}d(Ty,y)=\max_{h\in\mathcal{H}}\inf_{x\in X}(h(Tx)-h(x))$
where $\mathcal{H}$ is the set of metric functionals. Moreover if $d$ is bounded from below and if $\rho(T)>0$ then any function attaining the maximum is a horofunction.
In particular, any horofunction $h$ attaining the above maximum verifies $h(T^{k}x)-h(x)\geqslant k\rho(T)$ for all $k\in\mathbb{N}^{*}$, $x\in X$. Recently, Karlsson further refined this result, showing that when $T$ is an isometry, there exists a metric functional $h$ such that $h(Tx)-h(x)=\rho(T)$ for all $x\in X$, see [Karlsson2024].
1.2. Main results
In the present paper, we introduce a two-player game extension of the notion of escape rate. We consider a family of nonexpansive self-maps $(T_{ab})_{(a,b)\in\mathcal{A}\times\mathcal{B}}$ of a hemi-metric space $(X,d)$. We refer to the two players as Min and Max. The sets $\mathcal{A}$ and $\mathcal{B}$ are interpreted as the action spaces of Min and Max, respectively. We assume that Min and Max alternatively select actions $a_{1}\in\mathcal{A},b_{1}\in\mathcal{B},a_{2}\in\mathcal{A},b_{2}\in\mathcal{B},\dots$, each observing the previous actions of the other player before selecting their next action.
The escape rate induced by this infinite sequence of actions is defined as follows
$\limsup_{k\to\infty}\frac{d(T_{a_{k}b_{k}}\circ\dots\circ T_{a_{1}b_{1}}(\bar{x}),\bar{x})}{k}\enspace,$
where as above, the limit is independent of the choice of $\bar{x}$. Player Max aims to maximize this quantity, while Min aims to minimize it. We refer to this two-player zero-sum game as the Escape Rate Game. We always suppose that the action spaces are compact, and make a technical assumption which is satisfied in particular if the map $(a,b,x)\mapsto T_{ab}(x)$ is continuous (see 8 below). We are interested in the existence of uniform value, a notion studied by Mertens and Neyman [mertens_neyman]. Informally, a zero-sum repeated game is said to have a uniform value $\rho$ if both players can guarantee $\rho$ up to an arbitrary precision provided that they play for a sufficiently long time (see 17 below).
To solve this game, we introduce the following Shapley operator $S$, acting on the space $\operatorname{Lip}_{1}$ of Lipschitz functions of constant $1$ from $X$ to $(\mathbb{R},\delta_{1})$ with $\delta_{1}(x,y)=x-y$:
$Sv(x)\coloneqq\inf_{a\in\mathcal{A}}\sup_{b\in\mathcal{B}}v(T_{ab}(x))\enspace.$
Our first main result is the following maximin characterization of the value of the escape rate game.
Theorem 4.
The escape Rate Game has a uniform value $\rho$ given by
$\rho=\lim_{k\to\infty}\frac{[S^{k}d(\cdot,\bar{x})](\bar{x})}{k}$ $=\max\{\lambda\in\mathbb{R}\mid\exists\;v\in\operatorname{Lip}_{1},\;\lambda+v\leqslant Sv\}$
$=\max_{v\in\operatorname{Lip}_{1}}\inf_{x\in X}(Sv(x)-v(x))\enspace.$
We show in particular that player Max has an optimal stationary strategy, and that player Min has an optimal (non stationary) strategy.
We say that a function $f\in\operatorname{Lip}_{1}$ is distance-like if there exists a constant $\alpha$ such that $f(x)\geqslant d(x,\bar{x})+\alpha$. We denote by $\mathscr{D}$ the set of distance like functions. We say that a nonexpansive self map $T$ of a hemi metric space $(X,d)$ is an almost-isometry with constant $\gamma\in\mathbb{R}$ if it satisfies $d(T(x),T(y))\geqslant d(x,y)-\gamma$, for all $x,y\in X$. Having $S$ act on $\mathscr{D}$, we get the following dual minimax characterization of $\rho$.
Theorem 5.
Suppose that the maps $T_{ab}$ constitute a family of almost-isometries with a uniform constant. Then. the value of the escape rate game satisfies
$\rho=\inf\{\lambda\in\mathbb{R}\mid v\in\mathscr{D},\;\lambda+v\geqslant Sv\}=\inf_{v\in\mathscr{D}}\sup_{x\in X}(Sv(x)-v(x))\enspace.$
The infimum is not achieved in general. In contrast, in the maximin characterization, the supremum is always achieved, and the almost-isometry assumption is not needed.
The above minimax and maximin characterizations of the escape rate are reminiscent of “Collatz-Wielandt” type characterizations of the spectral radius, established by Nussbaum in the context of non-linear Perron-Frobenius theory, see [nussbaum86, nussbaumlemmens]. They may also be thought of as two player versions of the “Mañé-Conze-Guivarch lemma” characterizing the value of an ergodic control problem, see [Mane1992, Ma1996] and the discussion by Bousch in [Bousch2011]. The functions $v$ arising in these results are also closely related with the weak-KAM solutions defined by Fathi [fathi97b] for continuous time problems, see also [Fathi2004].
Our results apply in particular to the following example of hemi-metric. Suppose that $C\subset\mathbb{R}^{n}$ is a proper cone (i.e., convex, closed, pointed, of non-empty interior), and let $\leqslant_{C}$ denote the induced partial order on $\mathbb{R}^{n}$, so that $x\leqslant_{C}y$ if $y-x\in C$. The interior of $C$ can be equipped with the Funk hemi-metric, defined by
(1) $\operatorname{Funk}(x,y)=\log\inf\{\alpha>0\mid x\leqslant\alpha y\}\enspace.$
This hemi-metric is a key ingredient in Hilbert geometry [funk, Papadopoulos2008]. A self-map $T$ of the interior of $C$ is order preserving if $x\leqslant_{C}y$ implies $T(x)\leqslant_{C}T(y)$, for all $x,y$ in the interior of $C$, and it is positively homogeneous of degree $1$ if $T(\alpha x)=\alpha T(x)$ holds for all $\alpha>0$. These two properties are equivalent to $T$ being nonexpansive in the Funk hemi-metric. This allows us to define escape rate games for dynamics that are order preserving and positively homogeneous of degree $1$. Moreover, Corollary 33 below shows that for this special class of games, the condition that the maps $T_{ab}$ be almost-isometries in Theorem 5 can be dispensed with; to get a minimax characterization of the value, it suffices that each map $T_{ab}$ extend continuously to the closed cone $C$.
By specializing this approach to linear maps, we can solve matrix multiplication games, defined by Asarin et al. in [asarin_entropy]. In these games, one considers two compact sets of matrices $\mathcal{A}$ and $\mathcal{B}$ with entries in $\mathbb{C}$, and the players Min and Max alternatively select matrices $a_{1}\in\mathcal{A}$, $b_{1}\in\mathcal{B}$, $a_{2}\in\mathcal{A}$, …Player Max wants to maximize the payoff
(2) $\limsup_{k}\|a_{1}b_{1}\dots a_{k}b_{k}\|^{1/k}$
[…]
$\log\|a_{1}b_{1}\dots a_{k}b_{k}\|=\frac{1}{2}\operatorname{Funk}(T_{a_{k}b_{k}}\circ\dots\circ T_{a_{1}b_{1}}(I),I)$
where $I$ is the identity matrix.
By applying 4 and 5, we obtain the existence together with characterizations of the uniform value of matrix multiplication games with matrices in $\operatorname{GL}_{n}$. Similarly, 4 and Corollary 33 allow us to derive the same conclusion for matrices preserving the interior of a proper cone. In the case of nonnegative matrices, the existence of the value was previously known only under a restrictive “rectangularity condition” [asarin_entropy].
In the special case of “minimizer free” matrix multiplication games, in which Player Min is a dummy so that $\mathcal{A}=\{I\}$, the value of the escape rate game coincides with the joint spectral radius introduced by Rota and Strang [Rota1960]. When Player Max is a dummy, it coincides with the notion of lower spectral radius or joint spectral subradius introduced much later by Gurvits [Gurvits1995], and further studied by Jungers, Bochi and Morris [Jungers_2009, Bochi2014]. In particular, we show in 38, in the minimizer free case, that the existence of a distance-like function reaching the infimum in 5 is equivalent to the non-defectivity of the family of matrices.
Because of these special cases, we call competitive spectral radius the value of an escape rate game. It also includes as special one-player cases the notion of joint spectral radius of families of isometries introduced recently by Breuillard and Fujiwara [Breuillard2021], as well as the joint spectral radius of family of topical maps studied by Bousch and Mairesse [Bousch_2001]. We also note that Kozyakin introduced in [Kozyakin_2019] a notion of minimax joint spectral radius, which differs from ours (instead of alternating moves, one of the players first chooses an infinite sequence of matrices, leading to a Stackelberg game).
Concrete examples of competitive spectral radii arise in the study of topological entropies of transition systems [asarin_entropy] and in population dynamics, in particular in problems from mathematical biology and epidemiology, see in particular [billy, calvezgabriel, Chakrabarti2008]. They also arise in robust Markov decision processes, in which the second player represents the uncertainty, see [Goyal2023].
We illustrate our results by solving, in Section 6, a special class of escape rate game, in which the nonexpansive dynamics consists of translations acting on the Euclidean space. We give an explicit formula for a maximizing function $v$ in the maximin characterization of 5, which is based on a result of discrete geometry (Shapley-Folkman lemma).
2. Escape rate games: definition and first properties
2.1. Hemi-metric spaces
We shall need the following notion of asymmetric metric [GAUBERT_2011].
Definition 6 (Hemi-metric).
A map $d:X\times X\to\mathbb{R}$ is called a hemi-metric on the set $X$ if it satisfies the two following conditions:
  1. (1)
    $\forall(x,y,z)\in X^{3},\;d(x,z)\leqslant d(x,y)+d(y,z)$ (triangular inequality)
  2. (2)
    $\forall(x,y)\in X^{2},\;d(x,y)=d(y,x)=0$ if and only if $x=y$
A hemi-metric may take negative values. A basic example of hemi-metric on $\mathbb{R}^{n}$ is the map $\delta_{n}(x,y)\coloneqq\max_{i\in[n]}(x_{i}-y_{i})$.
Given a hemi-metric $d$, the map
$d^{\circ}(x,y)=\max(d(x,y),d(y,x))\enspace,$
is always an ordinary metric on $X$. We will refer to it as the symmetrized metric of $d$. In the sequel, we equip $X$ with the topology induced by $d^{\circ}$.
Given two hemi-metric spaces $(X,d)$ and $(Y,d^{\prime})$, we introduce the notion of a $1$-Lipschitz or nonexpansive function $v$ with respect to $d$ and $d^{\prime}$. Such a function verifies $d^{\prime}(v(x),v(y))\leqslant d(x,y)$ for all $x,y\in X$. We will especially be interested in the nonexpansive functions from $(X,d)$ to $(\mathbb{R},\delta_{1})$. It is easy to verify that a nonexpansive function between two hemi-metric spaces is nonexpansive with respect to the symmetrized metrics generated by their respective hemi-metrics. In particular, such a function is continuous with respect to the topologies induced by these symmmetrized metrics.
2.2. The Escape Rate Game
The Escape Rate Game is the following two-player deterministic perfect information game.
We fix two non-empty compact sets $\mathcal{A}$ and $\mathcal{B}$, that will represent the action spaces of the two players. We also fix a hemi-metric space $(X,d)$, which will play the role of the state space of the game. To each pair $(a,b)\in\mathcal{A}\times\mathcal{B}$, we associate a nonexpansive self-map $T_{ab}$ of $(X,d)$. The initial state $x_{0}\in X$ is given. We construct inductively a sequence of states $(x_{k})_{k\geqslant 0}$ as follows. Both players observe the current state $x_{k}$. At each turn $k\geqslant 1$, player Min chooses an action $a_{k}\in\mathcal{A}$ and then, after having observed the action $a_{k}$, player Max chooses an action $b_{k}\in\mathcal{B}$. The next state is given by $x_{k}=T_{a_{k}b_{k}}(x_{k-1})$.
First, we define the game in horizon $k$, denoted by $\Gamma_{k}$. There are $k$ successive turns, and player Min pays to Max the following amount:
(3) $J_{k}(a_{1}b_{1}\dots a_{k}b_{k})=d(T_{a_{k}b_{k}}\circ\dots\circ T_{a_{1}b_{1}}(x_{0}),x_{0})\enspace.$
Player Max aims to maximize this quantity, whereas player Min aims to minimize it.
We also define the escape rate game, denoted $\Gamma_{\infty}$. Here, an infinite number of turns are played, and player Min aims to minimize the escape rate
(4) $J_{\infty}(a_{1}b_{1}a_{2}b_{2}\dots)\coloneqq\limsup_{k\to\infty}\frac{J_{k}(a_{1}b_{1}\dots a_{k}b_{k})}{k}\enspace,$
while player Max wants to maximize it. It follows from the nonexpansiveness of the maps $T_{ab}$ that this limit is independent of the choice of $x_{0}$.
A strategy is a map that assigns a move to any finite history of the game (sequence of states and actions in successive turns). To formalize this notion, it will be convenient to use the notation $\mathcal{C}^{*}$ for the set of all finite sequences of elements from a set $\mathcal{C}$. Then, a strategy for Min is a map
$\sigma:(X\times\mathcal{A}\times\mathcal{B})^{*}\times X\to\mathcal{A},$
where the action chosen at turn $k+1$ depends on the history of states and actions
$(x_{0},a_{1},b_{1},x_{1},\ldots,a_{k},b_{k},x_{k}).$
Similarly, a strategy for Max is a map
$\tau:(X\times\mathcal{A}\times\mathcal{B})^{*}\times X\times\mathcal{A}\to\mathcal{B}\enspace.$
The set of strategies of Min (resp Max) will be written $\mathcal{A}^{s}$ (resp $\mathcal{B}^{s}$). A strategy $\sigma\in\mathcal{A}^{s}$ of Min is called stationary if it is independent of the turn $k$ and depends only on the position in the space, i.e. if for all $x\in X$, the restriction of $\sigma$, $\sigma_{x}:(X\times\mathcal{A}\times\mathcal{B})^{*}\times\{x\}\to\mathcal{A}$ is constant. Similarly, a strategy of Max is stationary if it is independent of the turn $k$, and depends only on the current state and last action of Min. Since the game is deterministic, any strategy can be rewritten as a function of the initial state and of the history of actions, but then the stationarity may not be visible.
For every pair $(\sigma,\tau)\in\mathcal{A}^{s}\times\mathcal{B}^{s}$, we denote by $J_{k}(\sigma,\tau)$ the payoff of $\Gamma_{k}$, as per (3), assuming that the sequence of actions $a_{1},b_{1},a_{2},b_{2},\dots$ is generated according to the strategies $\sigma$ and $\tau$, and similarly, we denote by $J_{\infty}(\sigma,\tau)$ the escape rate defined above.
Recall that the value of a game is the unique quantity that the players can both guarantee. Formally a game with strategy spaces $\mathcal{A}^{s}$ and $\mathcal{B}^{s}$ and payoff function $\mathcal{A}^{s}\times\mathcal{B}^{s}\to\mathbb{R}$, $(\sigma,\tau)\mapsto J(\sigma,\tau)$, has a value $\lambda\in\mathbb{R}$ if $\forall\epsilon>0,\;\exists\sigma^{*}\in\mathcal{A}^{s},\tau^{*}\in\mathcal{B}^{s}$ such that
$J(\sigma^{*},\tau)\leqslant\lambda+\epsilon\text{ and }J(\sigma,\tau^{*})\geqslant\lambda-\epsilon$
for all strategies $\sigma\in\mathcal{A}^{s}$ and $\tau\in\mathcal{B}^{s}$. Such strategies are called $\epsilon$-optimal strategies. When $\epsilon=0$, one gets optimal strategies, and in particular, $\lambda=J(\sigma^{*},\tau^{*})$. Equivalently, the game has a value if the following minimax and maximin coincide:
$\inf_{\sigma\in\mathcal{A}^{s}}\sup_{\tau\in\mathcal{B}^{s}}J(\sigma,\tau)=\sup_{\tau\in\mathcal{B}^{s}}\inf_{\sigma\in\mathcal{A}^{s}}J(\sigma,\tau)\enspace.$
The outer infimum and supremum are reached when the players have optimal strategies.
2.3. Examples of escape rate games
2.3.1. Vector Addition Games
[…]
$J_{\infty}(a_{1}b_{1}a_{2}b_{2}\dots)=\limsup_{k\to\infty}\frac{\|a_{1}+b_{2}+\dots+a_{k}+b_{k}\|}{k}\enspace.$
2.3.2. Conical Escape Rate Games
As in the introduction, we assume that $C\subset\mathbb{R}^{n}$ is a proper cone—i.e., a cone that is convex, closed, pointed, of non-empty interior. For all $y,z\in C\setminus\{0\}$, we define
$d(y,z)\coloneqq\operatorname{Funk}(y,z)\coloneqq\log\inf\{\alpha>0\mid y\leqslant\alpha z\}\enspace.$
In particular, $\operatorname{Funk}(y,z)=\log\inf\emptyset=+\infty$ if there is no $\alpha>0$ such that $y\leqslant\alpha z$. A part of the cone $C$ is an equivalence class of $C\setminus\{0\}$ for the relation $y\sim z$ if $\alpha^{\prime}z\leqslant y\leqslant\alpha z$ for some $\alpha,\alpha^{\prime}>0$. The map $\operatorname{Funk}(\cdot,\cdot)$ is referred to as the Funk metric. Its restriction to every part is a hemi-metric in the sense of Definition 6.
The symmetrized metric, known as Thompson’s part metric,
$d^{\circ}(y,z)=\operatorname{Th}(y,z)\coloneqq\max(\operatorname{Funk}(y,z),\operatorname{Funk}(z,y))\enspace,$
is a bona fide metric on every part of $C$. A related “metric” of interest is Hilbert’s projective metric
$\operatorname{Hil}(y,z)\coloneqq\operatorname{Funk}(y,z)+\operatorname{Funk}(z,y)\enspace.$
The term projective metric refers to the fact that $\operatorname{Hil}(y,z)=0$ iff $y$ and $z$ are proportional. See [nussbaum88, Papadopoulos2008, LN12] for background on these metrics.
For instance, if $C=(\mathbb{R}_{\geqslant 0})^{n}$, then $\operatorname{int}C=(\mathbb{R}_{>0})^{n}$, and for all $y,z\in(\mathbb{R}_{>0})^{n}$, setting $\operatorname{top}(x)\coloneqq\max_{i}x_{i}$,
$\operatorname{Funk}(y,z)=\log\max_{i\in[n]}\frac{y_{i}}{z_{i}}=\operatorname{top}(\log y-\log z),\qquad\operatorname{Th}(y,z)=\|\log y-\log z\|_{\infty}\enspace.$
Similarly, if $C=\mathcal{H}_{n}^{+}$ is the cone of hermitian positive semidefinite matrices, thought of as a real vector space, then, $\operatorname{int}C$ is the cone of positive definite hermitian matrices, and
$\operatorname{Funk}(y,z)=\log\lambda_{\max}(yz^{-1})\enspace.$
Remark 7.
Note that $\operatorname{Funk}(y,z)$ may take negative values. In contrast, Papadopoulos and Troyanov require in [Papadopoulos2008] weak metrics to be nonnegative, and so define the Funk metric to be $\max(0,\operatorname{Funk}(y,z))$. We work with the unsigned version as it captures more information.
We call order preserving a self-map $f$ of $\operatorname{int}C$ such that $x\leqslant_{C}y\implies f(x)\leqslant_{C}f(y)$. It is called positively homogeneous (of degree $1$) if for all $\lambda>0,x\in C$, $f(\lambda x)=\lambda f(x)$. Satisfying both these conditions is equivalent to the map being nonexpansive in the Funk metric. In particular, we denote by $\operatorname{End}(C)$ the set of linear self-maps of $\mathbb{R}^{n}$ which preserve the open cone $\operatorname{int}C$. Any element of $\operatorname{End}(C)$ is nonexpansive for the Funk metric.
For instance, if $C=(\mathbb{R}_{\geqslant 0})^{n}$, then, $\operatorname{End}(C)=\mathcal{M}^{n}_{+}$, the set of nonnegative matrices $M$ with at least one non-zero entry per column, acting on row vectors by $x\mapsto xM$. If $C=\mathcal{H}_{n}^{+}$, the set $\operatorname{End}(C)$ is harder to describe. It contains all completely positive maps of the form $T(x)=\sum_{i=1}^{k}a_{i}^{*}xa_{i},$ where $a_{1},\dots,a_{k}\in\mathbb{C}^{n\times n}$ are such that $\sum_{i=1}^{k}a_{i}^{*}a_{i}$ is invertible.
The conical escape rate game associated with the cone $C$ is defined as the escape rate game played on the interior of $C$, equipped with the hemi-metric $d(x,y)=\operatorname{Funk}(x,y)$. The two following examples are subclasses of such games.
2.3.3. Families of Nonnegative Matrices
We consider $C=(\mathbb{R}_{\geqslant 0})^{n}$. Let $\mathcal{A}$ and $\mathcal{B}$ be two non-empty compact subsets of +$\mathcal{M}^{n}_{+}$. Each pair of actions $(A,B)\in\mathcal{A}\times\mathcal{B}$ induces a self-map of $(\mathbb{R}_{>0})^{n}$, given by
$T_{AB}(x)\coloneqq xAB\enspace,$
where $x$ is interpreted as a row vector. Since $T_{AB}\in\operatorname{End}((\mathbb{R}_{\geqslant 0})^{n})$, it is nonexpansive in the Funk hemi-metric.
This yields an escape rate game, in which the payoff is given by
$\limsup_{k\to\infty}\frac{\operatorname{Funk}(x_{k},x_{0})}{k}$ $=\limsup_{k\to\infty}\log\max_{i\in\llbracket 1,n\rrbracket}(\frac{(x_{k})_{i}}{(x_{0})_{i}})^{\frac{1}{k}}$
$=\log(\limsup_{k\to\infty}(\max_{i\in\llbracket 1,n\rrbracket}(x_{k})_{i})^{1/k})$
$=\log(\limsup_{k\to\infty}\lVert x_{k}\rVert^{1/k}_{\infty})\enspace,$
where
$x_{k}=x_{0}A_{1}B_{1}\dots A_{k}B_{k}\enspace.$
Taking $x_{0}=(1,\dots,1)$ we get $\lVert x_{k}\rVert_{\infty}=\lVert A_{1}B_{1}\dots A_{k}B_{k}\rVert$, where for $A=(a_{ij})\in\mathbb{R}^{n\times n}$, $\lVert A\rVert=\max_{i}\sum_{k=1}^{n}\lvert a_{ik}\rvert$, and we recover (up to an exponential transformation) the payoff of a matrix multiplication game.
The reverse Funk metric
$\operatorname{RFunk}(x,y)\coloneqq\operatorname{Funk}(y,x)$
is also of interest. It yields an escape rate game in which the payoff is given by
$\limsup_{k\to\infty}\frac{\operatorname{RFunk}(x_{k},x_{0})}{k}$ $=\limsup_{k\to\infty}\log\max_{i\in\llbracket 1,n\rrbracket}(\frac{(x_{0})_{i}}{(x_{k})_{i}})^{1/k}$
$=-\liminf_{k\to\infty}\log\min_{i\in\llbracket 1,n\rrbracket}(\frac{(x_{k})_{i}}{(x_{0})_{i}})^{1/k}\enspace.$
Then, one player wants to maximize the smallest growth rate of the coordinates of $x^{k}$, whereas the other player wants to maximize it.
Such models arise in population dynamics, control of growth processes, and epidemiology: $x_{k}$ represents a population profile at time $k$; one player wishes to minimize the growth rate of the population whereas the other player wishes to maximize it, see [billy, calvezgabriel, Chakrabarti2008].
2.3.4. Families of Matrices in $\operatorname{GL}_{n}(\mathbb{C})$
Every matrix $M\in\operatorname{GL}_{n}(\mathbb{C})$ yields a congruence operator,
$\phi_{M}:X\mapsto M^{*}XM$
which is an automorphism of the cone of positive semidefinite Hermitian matrices, i.e., $\phi_{M}\in\operatorname{Aut}(\mathcal{H}_{n}^{+})$.
We say that a function $\nu$ on $\mathbb{R}^{n}$ is a symmetric Gauge function if it is invariant by permutation of the variables, convex and positively homogeneous of degree $1$. Every symmetric Gauge function $\nu$ yields a hemi-metric of Finsler type on $\operatorname{int}(\mathcal{H}_{n}^{+})$ ([nussbaumfinsler])
$d_{\nu}(X,Y)=\nu(\log(\operatorname{Spec}(XY^{-1})))\enspace.$
If $M\in\operatorname{GL}_{n}(\mathbb{C})$, then, we readily check that $\phi_{M}$ is an isometry of $(\operatorname{int}(\mathcal{H}_{n}^{+}),d_{\nu})$.
Consider $\mathcal{A},\mathcal{B}\subset\operatorname{GL}_{n}$ and the operators $(\phi_{AB})_{(A,B)\in\mathcal{A}\times\mathcal{B}}$, which are isometries, as $\phi_{AB}=\phi_{B}\circ\phi_{A}$ We denote by $\operatorname{Sing}$ the set of singular values of a matrix. We get that
$d_{\nu}(\phi_{A_{k}B_{A}}\circ\dots\circ\phi_{A_{1}B_{1}}(I),I)=2\nu(\log(\operatorname{Sing}(A_{1}B_{1}\dots A_{k}B_{k}))$
Therefore, depending on the Gauge function we can get some insights about the spectrum and singular values of the infinite product of matrices. In particular, when $\nu(x)=\operatorname{top}(x)=\max_{i}x_{i}$,
$d_{\operatorname{top}}(\phi_{A_{k}B_{A}}\circ\dots\circ\phi_{A_{1}B_{1}}(I),I)$ $=\operatorname{Funk}(\phi_{A_{k}B_{A}}\circ\dots\circ\phi_{A_{1}B_{1}}(I),I)$
$=2\log\|A_{1}B_{1}\dots A_{k}B_{k}\|$
[…]
3. Preliminary Results
To prove the existence and characterize the value of escape rate games, we use an operator approach (see [Rosenberg2001]). Throughout the paper, we make the following assumption.
Assumption 8.
For all $x\in X$, the maps $b\mapsto T_{ab}(x)$ and $a\mapsto T_{ab}(x)$ are continuous, and for all compact sets $K$, the set $\{T_{ab}(x)\ |\ (a,b,x)\in\mathcal{A}\times\mathcal{B}\times K\}$ is compact.
This is trivially verified if $(a,b,x)\mapsto T_{ab}(x)$ is continuous.
We denote by $\operatorname{Lip}_{1}$ the set of $1$-Lipschitz functions from the hemi-metric space $(X,d)$ to $(\mathbb{R},\delta_{1})$ (recall that $\delta_{1}(x,y)=x-y$). We define the following Shapley operator $S$, acting on $\operatorname{Lip}_{1}$,
(5) $Sv(x)\coloneqq\inf_{a\in\mathcal{A}}\sup_{b\in\mathcal{B}}v(T_{ab}(x))\enspace.$
The map $S$ is order preserving and commutes with the addition of a constant. Moreover, as proven in 10 below, 8 ensures that $S$ preserves the set $\operatorname{Lip}_{1}$ and is continuous for the topology of uniform convergence on compact sets. It also ensures that the infimum and supremum in (5) are attained.
Proposition 9.
Let $x_{0}\in X$, the set $\mathcal{L}_{x_{0}}=\{f\in\operatorname{Lip}_{1}\mid f(x_{0})=0\}$ is compact for the topology of compact convergence (otherwise called uniform convergence on compact sets).
Proof.
This set is obviously equi-continuous since it is equi-Lipschitz (for the hemi-metrics as well as the symmetrized metrics). Furthermore $\forall x\in X,\forall v\in\mathcal{L}_{x_{0}}$, we have $|v(x)|=|v(x)-v(x_{0})|\leqslant d^{\circ}(x,x_{0})$. So $\forall x$, $\{v(x)\mid v\in\mathcal{L}_{x_{0}}\}$ is a bounded set in $\mathbb{R}$ and therefore it is relatively compact. Now, by the Ascoli-Arzela theorem, we know that $\mathcal{L}_{x_{0}}$ is relatively compact for the topology of compact convergence (which coincides with the topology of pointwise convergence as $\mathcal{L}_{x_{0}}$ is equi-continuous [Kelley1975, Th 15 on p.232]). It is closed for this topology so it is compact.∎
Proposition 10.
Under 8, the operator $S$ preserves $\operatorname{Lip}_{1}$ and is continuous from $\operatorname{Lip}_{1}$ to itself for the topology of compact convergence.
Proof.
By 8, for any $x\in X$, and $a\in\mathcal{A}$, the map $b\in\mathcal{B}\mapsto T_{ab}(x)\in X$ is continuous. Let $v\in\operatorname{Lip}_{1}$. Since then $v$ is continuous from $X$ to $\mathbb{R}$ and $\mathcal{B}$ is compact, we get that $Sv(x)\leqslant\sup_{b\in\mathcal{B}}v(T_{ab}(x))<+\infty$. Similarly, $Sv(x)>-\infty$. This implies that $Sv(x)\in\mathbb{R}$ for all $x\in X$. By assumption the maps $T_{ab}$ are nonexpansive, all maps $x\mapsto v(T_{ab}(x))$ are in $\operatorname{Lip}_{1}$, so that $Sv$ also belongs to $\operatorname{Lip}_{1}$, as $\operatorname{Lip}_{1}$ is preserved by suprema and infima.
Let us now show that $S$ is continuous for the topology of compact convergence. Let $I$ be a directed set and consider a net $(v_{i})_{i\in I}$ converging to a function $v$ that belongs to $\operatorname{Lip}_{1}$. Let $K$ be a compact subset of $X$. We have
$\sup_{x\in K}|Sv_{i}(x)-Sv(x)|$ $\leqslant\sup_{(x,a,b)\in K\times\mathcal{A}\times\mathcal{B}}|v_{i}(T_{ab}(x))-v(T_{ab}(x))|$
$=\sup_{y\in Y}|v_{i}(y)-v(y)|$
where, by 8, $Y\coloneqq\{T_{ab}(x)|(x,a,b)\in K\times\mathcal{A}\times\mathcal{B}\}$ is compact. Therefore the RHS converges to 0 and so does the LHS. So $Sv_{i}\to Sv$ in the topology of compact convergence. ∎
Proposition 11.
There exists at least one additive eigenvalue of $S$ associated to an eigenvector in $\operatorname{Lip}_{1}$, i.e.
$\exists\lambda\in\mathbb{R}\text{ such that }\exists v\in\operatorname{Lip}_{1},Sv=\lambda+v$
Proof.
Let us recall that, given a topological space $X$, the space of continuous real-valued functions, $\mathcal{C}(X)$, equipped with the topology of uniform convergence on compact sets, which is defined by the family of semi-norms $\phi_{K}:f\mapsto\sup\{\lvert f(x)\rvert\mid x\in K\}$ where $K$ varies over the directed set of all compact subsets of $X$, is a locally convex topological vector space. Let us define $\tilde{S}v\coloneqq Sv-Sv(x_{0})$. It sends continuously $\mathcal{L}_{x_{0}}$ into itself. As $\mathcal{L}_{x_{0}}$, is a convex and compact subset of $\mathcal{C}(X)$, we can conclude that $\tilde{S}$ has a fixed point in $\mathcal{L}_{x_{0}}$ thanks to the generalization of the Schauder fixed point theorem by Tychonoff [Tychonoff_1935]. ∎
Note that this result becomes trivial if $d$ is a nonnegative hemi metric, since the null function would be an additive eigenvector with respect to the eigenvalue 00.
By the triangular inequality, the map $d(\cdot,x_{0}):x\mapsto d(x,x_{0})$ is in $\operatorname{Lip}_{1}$. We define
$s_{k}\coloneqq[S^{k}d(\cdot,x_{0})](x_{0})\enspace.$
Proposition 12.
We have
(6) $\rho\coloneqq\inf_{k\geqslant 1}\frac{s_{k}}{k}=\lim_{k}\frac{s_{k}}{k}\in\mathbb{R}\enspace.$
Proof.
We first show the following subadditivity property:
(7) $s_{k+l}\leqslant s_{k}+s_{l},\text{ for all }k,l\geqslant 1\enspace.$
Indeed, we have $S^{l}d(\cdot,x_{0}))\leqslant[S^{l}d(\cdot,x_{0})](x_{0})+d(\cdot,x_{0})=s_{l}+d(\cdot,x_{0})$ since $S^{l}d(\cdot,x_{0})$ is $1$-Lipschitz. Then, since $S^{k}$ is order preserving, and commutes with the addition of a constant, $S^{k+l}d(\cdot,x_{0})=S^{k}(S^{l}d(\cdot,x_{0}))\leqslant S^{k}(s_{l}+d(\cdot,x_{0}))=s_{l}+S^{k}(d(\cdot,x_{0}))$, and so, $[S^{k+l}d(\cdot,x_{0})](x_{0})\leqslant s_{l}+s_{k}$, showing (7). Then, it follows from Fekete subadditive lemma that the limit $\lim_{k}\frac{s_{k}}{k}$ does exists and that it coincides with $\rho\coloneqq\inf_{k}\frac{s_{k}}{k}$. In particular $\rho\leqslant s_{1}<+\infty$. It remains to show that $\rho>-\infty$.
Let $v=-d(x_{0},\cdot)$. By the triangular inequality, $v$ is in $\operatorname{Lip}_{1}$, so that $Sv(x_{0})\in\mathbb{R}$. We set
$\alpha\coloneqq[S(-d(x_{0},\cdot)](x_{0})=\inf_{a\in\mathcal{A}}\sup_{b\in\mathcal{B}}-d(x_{0},T_{ab}(x_{0}))\in\mathbb{R}\enspace.$
Again by the triangular inequality, we have $0=d(x_{0},x_{0})\leqslant d(x_{0},\cdot)+d(\cdot,x_{0})$. We deduce that
$s_{1}=[Sd(\cdot,x_{0})](x_{0})\geqslant[S(-d(x_{0},\cdot))](x_{0})=\alpha\enspace.$
We now show, by induction on $k$, that $s_{k}\geqslant\alpha k$ for all $k$. Since $S$ preserves $\operatorname{Lip}_{1}$, we have, for every $k\in\mathbb{N}^{*}$, $S^{k}d(\cdot,x_{0})\in\operatorname{Lip}_{1}$, and so
$[S^{k}d(\cdot,x_{0})](\cdot)\geqslant[S^{k}d(\cdot,x_{0})](x_{0})-d(x_{0},\cdot)=s_{k}-d(x_{0},\cdot)\enspace.$
Since $S$ is order preserving and commutes with the addition of a constant, we deduce $[S^{k+1}d(\cdot,x_{0})]\geqslant s_{k}+S(-d(x_{0},\cdot))$. Evaluating this inequality at point $x_{0}$, we get $s_{k+1}\geqslant s_{k}+\alpha$, which completes the induction.∎
The case in which there is only one player (Max) is of special interest. Then, the set $\mathcal{A}$ is trivial, being reduced to a single action $a$, and we set $T_{b}\coloneqq T_{ab}$ for all $b\in\mathcal{B}$. We call such games minimizer-free. The following proposition shows that in this case, the value of the escape rate game coincides with the asymptotic joint displacement introduced by Breuillard and Fujiwara [Breuillard2021, Claim 4]:
Proposition 13.
The value of a minimizer-free escape rate game coincides with the following quantity:
(8) $\rho_{\operatorname{disp}}=\lim_{k\to\infty}\inf_{x}\sup_{b_{1},\dots,b_{k}\in\mathcal{B}}\frac{d(T_{b_{k}}\circ\dots\circ T_{b_{1}}(x),x)}{k}\enspace.$
Proof.
In this special case, we have $s_{k}=\sup_{b_{1},\dots,b_{k}\in\mathcal{B}}d(T_{b_{k}}\circ\dots\circ T_{b_{1}}(x_{0}),x_{0})$, and $\rho=\lim_{k}s_{k}/k$. Moreover it is observed in [Breuillard2021, Claim 4] that for all choices of $x_{0}$, $\lim_{k}s_{k}/k=\rho_{\operatorname{disp}}$. ∎
Remark 14.
In our setting, at each turn, Min plays before Max, leading to a Shapley operator of the form $Sv(x)=\min_{a\in\mathcal{A}}\max_{b\in\mathcal{B}}v(T_{ab}(x))$. The variant in which Max plays first can be tackled by considering instead the operator $S^{\prime}v(x)=\max_{b\in\mathcal{B}}\min_{a\in\mathcal{A}}v(T_{ab}(x))$. The results which follow carry over to this case, with trivial modifications.
4. Maximin and Minimax Characterizations of the Value of Escape Rate Games
4.1. Existence and Maximin characterization of the value
The following theorem characterizes the quantity $\rho$ defined by (6) in terms of the non-linear sub-eigenvalues of $S$.
Theorem 15.
We have
$\rho=\max\{\lambda\in\mathbb{R}\mid\exists v\in\operatorname{Lip}_{1},\;\lambda+v\leqslant Sv\}\enspace.$
Proof.
We first observe that if $\lambda\in\mathbb{R}$ is such that $\exists v\in\operatorname{Lip}_{1}$, $\lambda+v\leqslant Sv$, then, $k\lambda+v\leqslant S^{k}v$, and since $d(\cdot,x_{0})\geqslant v-v(x_{0})$, $S^{k}d(\cdot,x_{0})\geqslant S^{k}(v-v(x_{0}))=S^{k}v-v(x_{0})$, and so $s_{k}/k\geqslant([S^{k}v](x_{0})-v(x_{0}))/k$, which entails that $\lim_{k}s_{k}/k\geqslant\lambda$. Therefore, $\rho\geqslant\sup\{\lambda\in\mathbb{R}\mid\exists v\in\operatorname{Lip}_{1},\;\lambda+v\leqslant Sv\}$.
To complete the proof, we need to construct a function $v$ such that $Sv\geqslant\rho+v$. Let $\lambda<\rho$. We first show the following result.
Lemma 16.
For any compact set $C\subset X$, $\exists k_{C}\in\mathbb{N}$ such that $S^{k_{C}}d(\cdot,x_{0})\geqslant k_{C}\lambda+d(\cdot,x_{0})$ on $C$.
Proof.
Let $C$ be a compact subset of $X$. As $d(\cdot,x_{0})\in\operatorname{Lip}_{1}$ and $S$ preserves $\operatorname{Lip}_{1}$, $\forall y\in C$, we have
$[S^{k}d(\cdot,x_{0})](y)$ $\geqslant[S^{k}d(\cdot,x_{0})](x_{0})-d(x_{0},y)$
$=s_{k}-d(x_{0},y)$
$=\lambda k+d(y,x_{0})+(s_{k}-\lambda k)-(d(y,x_{0})+d(x_{0},y))$
$\geqslant\lambda k+d(y,x_{0})+(s_{k}-\lambda k)-\max_{x\in C}(d(x,x_{0})+d(x_{0},x))$
Since $\rho-\lambda>0$ and $\rho=\lim_{k\to\infty}s_{k}/k$, there exists $k_{C}\in\mathbb{N}$ such that $(s_{k_{C}}-\lambda k_{C})-\max_{x\in C}(d(x,x_{0})+d(x_{0},x))\geqslant 0$, which gives the result.∎
We now finish the proof of 15. For convenience, we denote $d(\cdot,x_{0})$ by $w$. Let
(9) $\theta_{C}^{\lambda}\coloneqq\sup(w,Sw-\lambda,\dots,S^{k_{C}-1}w-(k_{C}-1)\lambda)\enspace.$
As the supremum of nonexpansive functions from $X$ to $\mathbb{R}$, the function $\theta_{C}^{\lambda}$ is also nonexpansive. Moreover, for all $x\in C$ we have
$[S\theta_{C}^{\lambda}](x)$ $\geqslant\sup([Sw](x),\dots,[S^{k_{C}}w](x)-(k_{C}-1)\lambda)$
$\geqslant\sup([Sw](x),\dots,[S^{k_{C}-1}w](x)-(k_{C}-2)\lambda,w(x)+\lambda)=\lambda+\theta_{C}^{\lambda}(x)\enspace.$
Let $\mathcal{K}$ denote the set of all compact subsets of $X$, ordered by inclusion, and consider the net $\{C\}_{C\in\mathcal{K}}$. After translating every function $\theta_{C}^{\lambda}$ by $-\theta_{C}^{\lambda}(x_{0})$, we assume that $\theta_{C}^{\lambda}(x_{0})=0$ holds for all $C\in\mathcal{K}$. Then, the net $(\theta_{C}^{\lambda})_{C\in\mathcal{K}}$ is included in $\mathcal{L}_{x_{0}}$, which is compact. Hence, there is a subnet $(\theta_{C}^{\lambda})_{C\in\mathcal{D}}$ along which $\theta_{C}^{\lambda}$ has a limit, denoted by $\theta^{\lambda}$, and this limit belongs to $\mathcal{L}_{x_{0}}$. Moreover, for all $x\in X$, the singleton $\{x\}$ belongs to $\mathcal{K}$, and then, by definition of a subnet, there is an element $C_{x}\in\mathcal{D}$ such that $\{x\}\subset C_{x}$. It follows that $[S\theta_{D}^{\lambda}](x)\geqslant\lambda+\theta_{D}^{\lambda}(x)$ holds for all $D\in\mathcal{D}$ such that $D\supset C_{x}$. Since the operator $S$ is continuous, taking the limit along the subnet, we deduce that $[S\theta^{\lambda}](x)\geqslant\lambda+\theta^{\lambda}(x)$ holds for all $x\in X$.
Therefore, for all $\lambda<\rho$ we have constructed a function $\theta^{\lambda}\in\mathcal{L}_{x_{0}}$ such that
$S\theta^{\lambda}\geqslant\lambda+\theta^{\lambda}\enspace.$
Using again the compactness of $\mathcal{L}_{x_{0}}$, we deduce that the net $\{\theta^{\lambda}\}_{\lambda<\rho}$, indexed by the set of real numbers smaller than $\rho$, admits a subnet converging to a function $\theta\in\mathcal{L}_{x_{0}}$. By continuity of $S$, this function verifies $S\theta\geqslant\rho+\theta$. Therefore, $\rho\leqslant\max\{\lambda\in\mathbb{R}\mid\exists v\in\operatorname{Lip}_{1},\;\lambda+v\leqslant Sv\}$. This shows the equality in 15. ∎
We will be interested with the existence of a uniform value as defined by Mertens and Neyman [mertens_neyman], which reinforces the notion of value of a game with mean-payoff objective.
Definition 17 (Uniform Value).
A zero-sum repeated game is said to have a uniform value $v_{\infty}$ if both players can guarantee $v_{\infty}$ up to an arbitrary precision provided that they play for a sufficiently long time. Formally $v_{\infty}$ is the uniform value of the game if for all $\epsilon>0$, there exist a pair of strategies $(\sigma_{\epsilon},\tau_{\epsilon})$ and a time $N$ such that for all $k\geqslant N$ and for all pairs of strategies $(\sigma,\tau)$,
$\frac{J_{k}(\sigma_{\epsilon},\tau)}{k}\leqslant v_{\infty}+\epsilon\qquad\qquad\frac{J_{k}(\sigma,\tau_{\epsilon})}{k}\geqslant v_{\infty}-\epsilon\enspace.$
This implies that the non zero-sum variant of the game, where Min still wishes to minimize $J(\sigma,\tau)$ but Max wishes to maximize
$J^{-}(\sigma,\tau)\coloneqq\liminf_{k\to\infty}J_{k}(a_{1}b_{1}\dots a_{k}b_{k})/k\enspace,$
instead of the limsup, still has the value $v_{\infty}$, meaning that:
$v_{\infty}\geqslant\inf_{\sigma}\sup_{\tau}\limsup_{k}J_{k}(\sigma,\tau)/k\geqslant\sup_{\tau}\inf_{\sigma}\liminf_{k}J_{k}(\sigma,\tau)/k\geqslant v_{\infty}\enspace.$
See [mertens_neyman] for more information on the uniform value; see also [bolte2013].
We also introduce the notion of $q$-cyclic strategy for player Min. In such a strategy, Min chooses a strategy for the game $\Gamma_{q}$ and repeats it indefinitely. Formally, we construct it inductively: at every $mq+1$ turn, $m\in\mathbb{N}$, player Min forgets the sequence of actions and states played during the $mq$ first turns and then plays the game $\Gamma_{q}$ as if it were starting from $x_{0}$ again. This means that for $m\in\mathbb{N},\ell=1,\dots,q$, the sequence of actions and states Min considers to choose their action at turn $mq+\ell$ is
$(x_{0},a_{mq+1},b_{mq+1},$ $T_{a_{mq+1}b_{mq+1}}(x_{0}),\dots$
$a_{mq+\ell-1},b_{mq+\ell-1},T_{a_{mq+\ell-1}b_{mq+\ell-1}}\circ\dots\circ T_{a_{mq+1}b_{mq+1}}(x_{0}))$
Note that the infinite sequence of actions played by Min in such a $q$-cyclic strategy will not be periodic unless player Max also selects actions in a cyclic way. We now give, and prove, a more detailed statement of 4
Theorem 18.
The escape rate game has a uniform value, which coincides with $\rho$. Furthermore, given a sub-eigenvector $v$ associated to $\rho$, player Max possesses a stationary optimal strategy obtained by maximizing $b\mapsto v(T_{a_{k+1}b}(x_{k}))$ at turn $k+1$. Player Min possesses a history dependent optimal strategy, and for all $\epsilon>0$, player Min admits an $\epsilon$-optimal strategy that is $q$-cyclic for some $q$.
Proof.
We first show the existence of the uniform value. We know from 15 that there exists $v\in\operatorname{Lip}_{1}$ such that $\rho+v\leqslant Sv$. Recall that any pair of strategies $(\sigma,\tau)$ of the two players Min and Max defines a sequence $(a_{k},b_{k})_{k\geqslant 0}$ of actions and that, starting from $x_{0}$, the sequence of states of the game is defined inductively by $x_{\ell}=T_{a_{\ell}b_{\ell}}(x_{\ell-1})$, for $\ell\geqslant 1$.
Let $\sigma$ be an arbitrary strategy of player Min. If $a_{1},b_{1},\dots,a_{k-1},b_{k-1}$ is the sequence of actions played up to stage $k-1$, the state $x_{k-1}$ of the game is given by
$x_{k-1}\coloneqq T_{a_{k-1}b_{k-1}}\circ\dots\circ T_{a_{1}b_{1}}(x_{0})\enspace,$
and $a_{k}$ is determined by $\sigma$ and the history. We then define the strategy $\tau^{*}$ of player Max, as follows: at step $k$, we select any action
$b_{k}\in\operatorname*{arg\,max}_{b\in\mathcal{B}}v(T_{a_{k}b}(x_{k-1}))\enspace.$
This is a stationary strategy as it depends only on the current state and the last action of Min. We get that the sequence of states generated by the strategies $\sigma$ and $\tau^{*}$ satisfies
$Sv(x_{k-1})$ $=\min_{a\in\mathcal{A}}\max_{b\in\mathcal{B}}v(T_{ab}(x_{k-1}))$
$\leqslant\max_{b\in\mathcal{B}}v(T_{a_{k}b}(x_{k-1}))=v(T_{a_{k}b_{k}}(x_{k-1}))=v(x_{k})\enspace.$
Since $\rho+v\leqslant Sv$, we deduce that $\rho+v(x_{k-1})\leqslant v(x_{k})$, for all $k\geqslant 1$, and so
$\rho\leqslant\frac{v(x_{k})-v(x_{0})}{k}\leqslant\frac{d(x_{k},x_{0})}{k}\enspace.$
It follows that for all $k$, and for all strategies $\sigma$ of Min, we have $\rho\leqslant J_{k}(\sigma,\tau^{*})$, showing that the second inequality required by the notion of uniform value, as per Definition 17, is satisfied for any $\epsilon>0$, with a choice of $\tau_{\epsilon}=\tau^{*}$ independent of $\epsilon$ and with $N=1$.
Now, to show the first inequality in Definition 17, we shall construct for all $\epsilon>0$ a $q$-cyclic $\epsilon$-optimal strategy for Min. Let $\epsilon>0$, as $\rho=\inf_{k\in\mathbb{N}^{*}}\frac{s_{k}}{k}$, there exists $q\in\mathbb{N}^{*}$ such that $\frac{s_{q}}{q}\leqslant\rho+\epsilon/2$, the value of the game in horizon $q$ exists and is equal to $s_{q}$. Let $\sigma_{q}$ be an optimal strategy of Min for the game $\Gamma_{q}$. We define $\sigma_{q}^{\infty}$ as the $q$-cyclic strategy of Min which consists of applying repeatedly $\sigma_{q}$. Let $\tau\in\mathcal{B}^{s}$. As usual, we denote $(a_{i},b_{i})_{i\in\mathbb{N}^{*}}$ the sequence of actions induced by the strategies $\sigma_{q}^{\infty}$ and $\tau$ and for all $m\in\mathbb{N}$, we define the operator played from turn $(m-1)q+1$ to turn $mq$, $\widetilde{T}_{w_{m}}\coloneqq T_{a_{mq}b_{mq}}\circ\dots\circ T_{a_{(m-1)q+1}b_{(m-1)q+1}}$. As $\sigma_{q}$ is an optimal strategy for the game $\Gamma_{q}$ we have for all $m\in\mathbb{N}$
$d(\widetilde{T}_{w_{m}}(x_{0}),x_{0})\leqslant s_{q}$
So using the properties of $d$ and the nonexpansiveness of the operators we get
$J_{mq}(\sigma_{q}^{\infty},\tau)$ $=d(\widetilde{T}_{w_{m}}\circ\dots\circ\widetilde{T}_{w_{1}}(x_{0}),x_{0})$
$\leqslant d(\widetilde{T}_{w_{m}}\circ\dots\circ\widetilde{T}_{w_{1}}(x_{0}),\widetilde{T}_{w_{m}}(x_{0}))+d(\widetilde{T}_{w_{m}}(x_{0}),x_{0})$
$\leqslant d(\widetilde{T}_{w_{m-1}}\circ\dots\circ\widetilde{T}_{w_{1}}(x_{0}),x_{0})+d(\widetilde{T}_{w_{m}}(x_{0}),x_{0})$
[…]
$\leqslant d(\widetilde{T}_{w_{m}}(x_{0}),x_{0})+\dots+d(\widetilde{T}_{w_{1}}(x_{0}),x_{0})\leqslant ms_{q}\leqslant mq(\rho+\epsilon/2)\enspace.$
We set
(10) $M\coloneqq\max_{(a,b)\in A\times B}d(T_{ab}(x_{0}),x_{0})\enspace,$
which is finite by 8.
By the same arguments as above, we have for $0\leqslant\ell<q$, and $k=mq+\ell$,
$J_{k}(\sigma_{q}^{\infty},\tau)$ $\leqslant d(T_{a_{mq+\ell}b_{mq+\ell}}\circ\dots\circ T_{a_{mq+1}b_{mq+1}}(x_{0}),x_{0})+ms_{q}$
$\leqslant d(T_{a_{mq+\ell}b_{mq+\ell}}(x_{0}),x_{0})+\dots+d(T_{a_{mq+1}b_{mq+1}}(x_{0}),x_{0})+ms_{q}$
$\leqslant\ell M+ms_{q}\leqslant\ell M+mq(\rho+\frac{\epsilon}{2})$
(11) $=\ell(M-\rho-\frac{\epsilon}{2})+k(\rho+\frac{\epsilon}{2})\leqslant(q-1)(M^{+}-\rho^{-}-\frac{\epsilon}{2})+k(\rho+\frac{\epsilon}{2})\enspace,$
where $a^{+}\coloneqq\max(a,0)$ and $a^{-}\coloneqq\min(a,0)$. Taking $N$ such that $(q-1)(M^{+}-\rho^{-}-\epsilon/2)/N\leqslant\epsilon/2$, we deduce that for all $k\geqslant\max(q,N)$, and for all strategies $\tau$ of Max, $J_{k}(\sigma_{q}^{\infty},\tau)/k\leqslant\rho+\epsilon$, completing the proof that $\rho$ is the uniform value of the game. We also showed that Max possesses a stationary optimal strategy, whereas Min possesses cyclic $\epsilon$-optimal strategies for all $\epsilon>0$.
Now, to construct an optimal strategy for Min, we follow the construction of a cyclic strategy but instead of keeping a fixed period, we increase it by $1$ every time. This means that for the first turn Min plays an optimal strategy associated with $\Gamma_{1}$. Then, he forgets the actions played so far (the other player can remember them) and plays an optimal strategy for the game $\Gamma_{2}$. Then, he forgets the first turn, and for the $2$ following turns, plays an optimal strategy associated with $\Gamma_{2}$. He repeats this operation, i.e., after the $1+\dots+q$ first turns, he forgets the past and plays an optimal strategy associated with $\Gamma_{q+1}$. We denote by $\sigma^{*}$ this strategy of Min.
By the same arguments as above (see the derivation of (11)), we obtain for every $q\in\mathbb{N}^{*}$, $0\leqslant\ell\leqslant q$, and $\tau\in\mathcal{B}^{s}$ the following property:
$J_{1+\dots+q+\ell}(\sigma^{*},\tau)\leqslant\ell M+s_{1}+\dots+s_{q}\enspace.$
Dividing by $1+\dots+q+\ell$, we deduce, that for all $q\geqslant 1$,
$\frac{J_{1+\dots+q+\ell}(\sigma^{*},\tau)}{1+\dots+q+\ell}\leqslant\frac{qM^{+}+s_{1}+\dots+s_{q}}{1+\dots+q}=\frac{2M^{+}}{q+1}+\frac{s_{1}+\dots+s_{q}}{1+\dots+q}\enspace.$
Observe that any integer $k\geqslant 1$ can be written in a unique way as $1+2+\dots+q+\ell$ with $q\geqslant 1$ and $0\leqslant\ell\leqslant q$. We denote by $q_{k}$ the unique integer $q$ arising in this decomposition. Then, for all $k\in\mathbb{N}^{*}$ and $\tau\in\mathcal{B}^{s}$,
$\frac{J_{k}(\sigma^{*},\tau)}{k}\leqslant\frac{2M^{+}}{q_{k}+1}+\frac{s_{1}+\dots+s_{q_{k}}}{1+\dots+q_{k}}$
As $k$ goes to +$\infty$+ ∞, so does $q_{k}$. Moreover, since $s_{k}/k$ tends to $\rho$ as $k$ tends to infinity, the weighted Cesaro sum $(s_{1}+\dots+s_{k})/(1+2+\dots+k)$ tends to the same limit (see Theorem 43, p100 of [hardy]). Then, for all $\epsilon>0$, we can find an integer $N$ such that for all $k\geqslant N$, $\frac{s_{1}+\dots+s_{q_{k}}}{1+\dots+q_{k}}\leqslant\rho+\epsilon/2$, and $\frac{2M^{+}}{q_{k}+1}\leqslant\epsilon/2$, so that $J_{k}(\sigma^{*},\tau)/k\leqslant\rho+\epsilon$, showing that $\sigma^{*}$ is an optimal policy of Min in the escape rate game. ∎
We have the direct corollaries for vector addition games, conical games, and matrix multiplication games.
Corollary 19.
Vector addition games have a uniform value.
Corollary 20.
Conical escape rate games have a uniform value. In particular, matrix multiplication games with matrices that are either invertible or preserve the interior of a proper cone have a uniform value.
4.2. Distance-like functions and minimax characterization of the value
Recall that we say a $1$-Lipschitz function $v:X\to\mathbb{R}$ is distance-like if, for some $x_{0}\in X$, there exists a constant $\alpha\in\mathbb{R}$ such that
$v(x)\geqslant\alpha+d(x,x_{0}),\quad\forall x\in X\enspace.$
Distance-like functions allow one to bound from above the value of an escape rate game.
Proposition 21.
Suppose that there exists a distance-like function $v$ such that
$Sv\leqslant\lambda+v\enspace,$
for some $\lambda\in\mathbb{R}$. Then, the value $\rho$ of the escape rate game satisfies $\rho\leqslant\lambda$. Moreover, by selecting at state $x$ an action $a\in\operatorname*{arg\,min}_{a\in\mathcal{A}}\sup_{b\in\mathcal{B}}v(T_{ab}(x))$, we obtain a stationary strategy for player Min which guarantees $\lambda$.
Proof.
By using this strategy, we get, by a immediate induction, that for any sequence of actions $(b_{k})_{k\in\mathbb{N}^{*}}$ played by Max, setting $x_{k}=T_{a_{k}b_{k}}\circ\dots\circ T_{a_{1}b_{1}}(x_{0})$,
$d(x_{k},x_{0})+\alpha$ $\leqslant v(x_{k})=v(T_{a_{k}b_{k}}(x_{k-1}))\leqslant\sup_{b\in\mathcal{B}}v(T_{a_{k}b}(x_{k-1}))=Sv(x_{k-1})$
$\leqslant v(x_{k-1})+\lambda\leqslant\cdots\leqslant v(x_{0})+k\lambda\enspace,$
showing that player Min can guarantee $\lambda$. Hence, the value of the escape rate game satisfies $\rho\leqslant\lambda$. ∎
Coupling 21 with 4 we get the following result.
Corollary 22.
Suppose that there exists a distance-like function $v$ and $\lambda\in\mathbb{R}$ such that
$Sv=\lambda+v\enspace,$
then, $\lambda$ is the value of the game.
We shall say that a nonexpansive map $T:X\to X$ is an almost-isometry if for some $\gamma\geqslant 0$,
(12) $d(T(x),T(y))\geqslant d(x,y)-\gamma\enspace,\forall x,y\in X\enspace.$
In particular, an isometry is an almost-isometry (with $\gamma=0$). We shall say that the nonexpansive maps $T_{ab}$ with $(a,b)\in A\times B$ constitute a uniform family of almost-isometries if every map $T=T_{ab}$ satisfies (12) for some constant $\gamma\geqslant 0$ independent of the choice of $(a,b)\in A\times B$.
Proposition 23.
Suppose that the maps $T_{ab}$ with $(a,b)\in A\times B$ constitute a uniform family of almost-isometries. Then, the Shapley operator $S$ preserve the space of distance-like functions $\mathscr{D}$.
Proof.
Since $S$ commutes with the addition of a constant, it suffices to show that $Sd(\cdot,x_{0})$ is distance like. Then, defining $M$ as per (10), we get that for all $(a,b)\in A\times B$ and $x\in X$,
$d(T_{ab}(x),x_{0})\geqslant d(T_{ab}(x),T_{ab}(x_{0}))-d(x_{0},T_{ab}(x_{0}))\geqslant d(x,x_{0})-\gamma-M\enspace.$
This implies that $Sd(\cdot,x_{0})\geqslant d(\cdot,x_{0})-\gamma-M$. ∎
We now prove the minimax characterization of the value of the game, stated as 5 in the introduction.
Proof of 5.
It follows from 21 that $\rho\leqslant\inf\{\lambda\in\mathbb{R}\mid v\in\mathscr{D},Sv\leqslant\lambda+v\}$. The proof of the reverse inequality relies on the following lemma.
Lemma 24.
For all $\lambda>\rho$ there exists $k\in\mathbb{N}^{*}$ such that $S^{k}d(\cdot,x_{0})\leqslant k\lambda+d(\cdot,x_{0})$.
Proof.
We have $S^{k}d(\cdot,x_{0})\in\operatorname{Lip}_{1}$ so
$S^{k}d(\cdot,x_{0})\leqslant[S^{k}d(\cdot,x_{0})](x_{0})+d(\cdot,x_{0})=s_{k}+d(\cdot,x_{0})\enspace.$
Since $\inf_{k\geqslant 1}s_{k}/k=\rho$ and $\lambda-\rho>0$, there exists $k\geqslant 1$ such that $s_{k}/k\leqslant\lambda$. Then, $Sd(\cdot,x_{0})\leqslant k\lambda+d(\cdot,x_{0})$. ∎
The end of the proof relies on a dual version of the “sup-averaging” argument (9) used in the proof of 15. We take $k$ as in (24), and define
$\theta\coloneqq\inf\{d(\cdot,x_{0}),Sd(\cdot,x_{0})-\lambda,\dots,S^{k-1}d(\cdot,x_{0})-(k-1)\lambda\}\enspace.$
Since, by 23, $S$ preserves the set of distance-like functions, and since this set is stable by taking pointwise infima of finite families of elements of this set,, the function $\theta$ is distance-like. Moreover, since $S$ is order preserving, we have
$S\theta$ $\leqslant\inf\{Sd(\cdot,x_{0}),S^{2}d(\cdot,x_{0})-\lambda,\dots,S^{k}d(\cdot,x_{0})-(k-1)\lambda\}$
$\leqslant\inf\{Sd(\cdot,x_{0}),S^{2}d(\cdot,x_{0})-\lambda,\dots,S^{k-1}d(\cdot,x_{0})-(k-2)\lambda,d(\cdot,x_{0})+\lambda\}$
$=\theta+\lambda\qed$
The infimum in 5 is not reached in all generality, i.e. for some families of isometries there cannot exist a distance-like function such that $Sv\leqslant\rho+v$ as shown by the following example.
Example 25.
Take $X$ as being the Poincaré half space $\{z\in\mathbb{C}\mid\Im(z)>0\}$ equipped with the Poincaré Metric
$d(z,z^{\prime})=2\operatorname{arcsinh}(\frac{\lVert z-z^{\prime}\rVert}{\sqrt{\Im(z)}\sqrt{\Im(z^{\prime})}})\enspace,$
where $\Im(\cdot)$ denotes the imaginary part of a complex number. We suppose the two players are dummies, so that a single dynamics is available $T(z)=z+1$. It is trivial to see that $\lim_{k\to\infty}\frac{d(T^{k}(i),i)}{k}=0$. Then, if there exists a distance-like function $v$ (of constant $\gamma\in\mathbb{R}$ and center $x_{0}$) such that for all $x\in X$, $v(Tx)\leqslant v(x)$, we would have for all $k\in\mathbb{N}^{*}$
$\gamma+2\frac{\operatorname{arcsinh}(k)}{\Im(x_{0})}\leqslant v(T^{k}(x_{0}))\leqslant v(x_{0})$
which is absurd as $\lim_{k\to\infty}\operatorname{arcsinh}(k)=+\infty$.
Combining 21 and 5, we obtain:
Corollary 26.
Suppose that the maps $T_{ab}$ with $(a,b)\in A\times B$ constitute a uniform family of almost-isometries. Then, for all $\lambda>\rho$, player Min has a stationary strategy which guarantees $\lambda$. ∎
5. Conical escape rate games
We now further investigate conical escape rate games, introduced in Section 2.3.2. We consider a family of order-preserving and positively homogeneous maps $(T_{ab})_{a\in\mathcal{A},b\in\mathcal{B}}$ acting on the interior of a proper cone $C\subset\mathbb{R}^{n}$.
Recall that every map $T_{ab}$ that is order preserving and positively homogeneous of degree $1$ is nonexpansive on the space $X=\operatorname{Int}C$ equipped with the hemi-metric $d(x,y)=\operatorname{Funk}(x,y)$. By Theorem 4, the associated escape rate game has a uniform value $\rho$, which admits a maximin characterization. In contrast, the minimax characterization (5) requires a uniform family of almost isometries. We next show that, owing to the particular conical setting, the almost-isometry assumption can be dispensed with.
We consider a cross section of $C$ of the form $\Delta\coloneqq\{x\in C\mid\langle x,e^{*}\rangle=1\}$ where $e^{*}$ is a point in the interior of the dual cone $C^{*}$. We recall the following result.
Lemma 27 (Lemma 1.2.4 of [nussbaumlemmens]).
The cross section $\Delta$ of $C$ is compact for the Euclidean topology.
[…]
Lemma 28.
There exists constants $\beta,\beta^{\prime}\in\mathbb{R}$ such that for all $x\in C\setminus\{0\}$,
(13) $-\beta+\log\langle x,e^{*}\rangle$ $\leqslant\operatorname{Funk}(x,x_{0})\leqslant\beta+\log\langle x,e^{*}\rangle,$
(14) $-\beta^{\prime}+\log\|x\|$ $\leqslant\operatorname{Funk}(x,x_{0})\leqslant\beta^{\prime}+\log\|x\|\enspace.$
Proof.
Let $\Delta^{*}\coloneqq\{\varphi\in C^{*}\mid\varphi(x_{0})=1\}$. Applying Lemma 27 to the dual cone $C^{*}$, we get that $\Delta^{*}$ is compact. We have
$\operatorname{Funk}(x,x_{0})=\log\max_{\varphi\in\Delta^{*}}\varphi(x)\enspace,$
the maximum being achieved, see e.g. [GAUBERT_2011, Lemma 27]. Since the map $(x,\varphi)\mapsto\varphi(x)$ is continuous, and $\Delta^{*}$ is compact, it follows that $\operatorname{Funk}(\cdot,x_{0})$ is continuous, and so, it is bounded on $\Delta$, so that there exists a constant $\beta\in\mathbb{R}$ such that $-\beta\leqslant\operatorname{Funk}(x,x_{0})\leqslant\beta$ for all $x\in\Delta$. By homogeneity, we have $\operatorname{Funk}(\lambda x,x_{0})=\operatorname{Funk}(x,x_{0})+\log(\lambda)$ for all $\lambda>0$, and the same is true of $\log\langle x,e^{*}\rangle$, which shows that (13) holds. The proof of (14) is similar. ∎
[…]
Assumption 29.
For all $(a,b)\in\mathcal{A}\times\mathcal{B},T_{ab}$ can be continuously extended to $C$ with respect to the norm topology of $C$. Furthermore, the map $\mathcal{A}\times\mathcal{B}\times C\to C,(a,b,x)\mapsto T_{ab}(x)$ is continuous, $C$ being equipped with this topology.
This assumption is trivially verified if the maps $T_{ab}$ are linear, and in particular in Examples 2.3.3 and 2.3.4. Note also that if the cone $C$ is polyhedral, any self-map of the interior of $C$ that is order preserving and positively homogeneous of degree $1$ can be extended continuously to $C$, see [MR1960046].
We now reduce the conical escape rate game played on the state space $\operatorname{int}C$ to a new game $\Gamma_{\Delta}$ played on the state space $\Delta$. The two players Min and Max still alternatively choose actions in $\mathcal{A}$ and $\mathcal{B}$. The state of the game now evolves according to the normalized operators $\hat{T}_{ab}(x)=\frac{T_{ab}(x)}{\langle T_{ab}(x),e^{*}\rangle}$. We suppose at each turn $k$, Min gives to Max  a payment equal to $\log(\langle T_{a_{k}b_{k}}(\hat{x}_{k-1}),e^{*}\rangle)$, where $\hat{x_{k}}\coloneqq\hat{T}_{a_{k}b_{k}}\circ\dots\circ\hat{T}_{a_{1}b_{1}}(x_{0})$. So, the payoff of the game in horizon $k$ is
$\sum_{k=1}^{n}\log(\langle T_{a_{k}b_{k}}(\hat{x}_{k-1}),e^{*}\rangle)=\log(\langle T_{a_{n}b_{n}}\circ\dots\circ T_{a_{1}b_{1}}(x_{0}),e^{*}\rangle)\enspace.$
The payoff of an infinite play of the game $\Gamma_{\Delta}$ is given by
$\limsup_{k\to\infty}\frac{\log(\langle T_{a_{k}b_{k}}\circ\dots\circ T_{a_{1}b_{1}}(x_{0}),e^{*}\rangle)}{k}$
By (13), this is identical to the payoff $\limsup_{k\to\infty}\operatorname{Funk}(x_{k},x_{0})/k$, so that the conical escape rate game is equivalent to the game $\Gamma_{\Delta}$ starting from any state $x_{0}$ in the relative interior of $\Delta$.
We now introduce $F$, the Shapley operator of $\Gamma_{\Delta}$, acting on the space of continuous functions on $\Delta$, $\mathscr{C}(\Delta)$, as follows:
$Fv(x)\coloneqq\inf_{a\in\mathcal{A}}\sup_{b\in\mathcal{B}}\Big{\{}\log(\langle T_{ab}(x),e^{*}\rangle)+v\big{(}\frac{T_{ab}(x)}{\langle T_{ab}(x),e^{*}\rangle}\big{)}\Big{\}}$
We consider the subspace $\operatorname{CLip}_{1}(\Delta)$ of $\mathscr{C}(\Delta)$, consisting of continuous functions $v:\Delta\to\mathbb{R}$ such that $v(x)-v(y)\leqslant\operatorname{Funk}(x,y)$ as soon as $\operatorname{Funk}(x,y)<\infty$. Observe that if $T:C\to\mathbb{R}$ is continuous, order preserving and positively homogeneous of degree $1$, then the restriction of $T$ to $\Delta$ belongs to $\operatorname{CLip}_{1}(\Delta)$.
Lemma 30.
The map $F$ preserves $\operatorname{CLip}_{1}(\Delta)$.
Proof.
Let $v\in\operatorname{CLip}_{1}(\Delta)$. By assumption $(x,a,b)\mapsto(T_{ab}(x))$ is continuous so is
$\phi:(x,a,b)\mapsto\log(\langle T_{ab}(x),e^{*}\rangle)+v\big{(}\frac{T_{ab}(x)}{\langle T_{ab}(x),e^{*}\rangle}\big{)}$
If $f:X\times Y\to\mathbb{R}$ is a continuous function and $Y$ is compact then $x\mapsto\sup_{y\in Y}f(x,y)$, and $x\mapsto\inf_{y\in Y}f(x,y)$ are both continuous function. Then as $\mathcal{B}$ is compact, $(x,a)\mapsto\sup_{b\in\mathcal{B}}\phi(x,a,b)$ is continuous and as $\mathcal{A}$ is compact, $Fv$ is continuous.
Let $x,y\in\Delta$ be such that $\operatorname{Funk}(x,y)<\infty$. Then, for all $(a,b)\in\mathcal{A}\times\mathcal{B}$, we have
$\phi(x,a,b)-\phi(y,a,b)$
$=\log(\langle T_{ab}(x),e^{*}\rangle)-\log(\langle T_{ab}(y),e^{*}\rangle)+v\big{(}\frac{T_{ab}(x)}{\langle T_{ab}(x),e^{*}\rangle}\big{)}-v\big{(}\frac{T_{ab}(y)}{\langle T_{ab}(y),e^{*}\rangle}\big{)}$
$\leqslant\log(\langle T_{ab}(x),e^{*}\rangle)-\log(\langle T_{ab}(y),e^{*}\rangle)+\operatorname{Funk}\big{(}\frac{T_{ab}(x)}{\langle T_{ab}(x),e^{*}\rangle},\frac{T_{ab}(y)}{\langle T_{ab}(y),e^{*}\rangle}\big{)}$
$=\operatorname{Funk}(T_{ab}(x),T_{ab}(y))\leqslant\operatorname{Funk}(x,y)$
We deduce that $[Fv](x)-[Fv](y)\leqslant\operatorname{Funk}(x,y)$. ∎
We now recall the following definition from [GAUBERT_2011]. It captures a mild form of nonpositive curvature in the sense of Busemann.
Definition 31.
We say that the hemi-metric space $(X,d)$ is metrically star-shaped with center $x^{*}$ if there exists a family of geodesics $(\gamma_{y})_{y\in X}$, such that $\gamma_{y}$ joins $x^{*}$ to $y$ and such that the following inequality is satisfied for every $(y,z)\in X^{2}$ and $s\in[0,1]$:
$d(\gamma_{y}(s),\gamma_{z}(s))\leqslant sd(y,z)$
We also require that for every $y$, $\lim_{s\to 1}d(y,\gamma_{y}(s))=0$.
We now consider the space $X=\operatorname{CLip}_{1}(\Delta)$, equipped with the hemi-metric $\delta:(f,g)\mapsto\max_{x\in\Delta}\big{(}f(x)-g(x)\big{)}$. We take the zero function as the center, together with the choice of geodesic $\gamma_{f}$ from the zero function to a function $f\in\operatorname{CLip}_{1}(\Delta)$, such that $\gamma_{f}(s)=sf$, for all $s\in[0,1]$. Then, the space $\operatorname{CLip}_{1}(\Delta)$ is metrically star shaped. Moreover, the symmetrized metric of $\delta$ is the sup-norm on $\operatorname{CLip}_{1}(\Delta)$.
Theorem 32.
The value $\rho$ of a conical escape game satisfies
(15) $\rho$ $=\lim_{k\to\infty}\frac{\delta(F^{k}(0),0)}{k}$
(16) $\ =\inf_{v\in\operatorname{CLip}_{1}(\Delta)}\max_{x\in\Delta}\big{\{}[F(v)](x)-v(x)\big{\}}$
(17) $=\inf\{\lambda\in\mathbb{R}\mid\exists v\in\operatorname{CLip}_{1}(\Delta),Fv\leqslant\lambda+v\}\enspace.$
Proof.
We already observed that the existence of the (uniform) value $\rho$ follows from Theorem 4.
Since the space $\operatorname{CLip}_{1}(\Delta)$ is metrically star shaped and complete, Theorem 3 implies that the three expressions at the right-hand sides of (15)–(17) have the same value. Let $\bar{\rho}$ denote this common value. Let $\lambda>\bar{\rho}$, and let $v\in\operatorname{CLip}_{1}(\Delta)$ be such that $Fv\leqslant\lambda+v$. Let us consider the stationary strategy $\sigma^{*}$ of Min, which in state $x\in C\setminus\{0\}$, selects any action $a\in\mathcal{A}$ that achieves the minimum in the expression of $[Fv](x)$. Then, applying any strategy $\tau$ of Max, we get a trajectory $x_{0},x_{1},\dots$ such that
$\log\langle x_{k},e^{*}\rangle+v(\frac{x_{k}}{\langle x_{k},e^{*}\rangle})\leqslant k\lambda+v(\frac{x_{0}}{\langle x_{0},e^{*}\rangle})$
By 28, $\operatorname{Funk}(x_{k},x_{0})\leqslant\beta+\log\langle x_{k},e^{*}\rangle$. Since $v$ is bounded by some constant $M$, it follows that $\operatorname{Funk}(x_{k},x_{0})\leqslant\beta+M+k\lambda$, which shows that Min can guarantee any $\lambda>\bar{\rho}$.
We next show that Max can guarantee $\bar{\rho}$, using Corollary 23 from [GAUBERT_2011]. It shows that there exists $x^{*}\in\Delta$ (not necessarily in the interior of $C$) such that for all $k\in\mathbb{N}$
$\frac{F^{k}(0)(x^{*})}{k}\geqslant\bar{\rho}\enspace.$
We now construct the stationary strategy $\tau^{*}$, such that, at stage $k$, assuming that Min just played action $a_{k}$, Max plays any action $b_{k}$ achieving the maximum in
$\sup_{b\in\mathcal{B}}\left(\log\langle T_{a_{k}b}(x^{*}),e^{*}\rangle+[F^{k-1}(0)](\frac{T_{a_{k}b}(x^{*})}{\langle T_{a_{k}b}(x^{*}),e^{*}\rangle})\right)\enspace.$
For all strategies $\sigma$ of Min, we get $\log\langle T_{a_{k}b_{k}}\circ\dots\circ T_{a_{1}b_{1}}(x^{*}),e^{*}\rangle\geqslant k\bar{\rho}$. Since $x_{0}$ belongs to the interior of $C$, there exists a constant $\alpha>0$ such that $x_{0}\geqslant\alpha x^{*}$. We deduce that $\log\langle T_{a_{k}b_{k}}\circ\dots\circ T_{a_{1}b_{1}}(x_{0}),e^{*}\rangle\geqslant k\bar{\rho}-\log\alpha$ which entails, by 28, that Max can guarantee $\bar{\rho}$ by playing $\tau^{*}$. So $\bar{\rho}$ coincides with the value $\rho$ of the escape rate game. ∎
Let us denote by $\mathscr{D}_{C}$ the space of continuous functions (for the euclidean topology) on $C$ which are distance-like (for the Funk metric) on the interior of $C$. For a function $v\in\operatorname{CLip}_{1}(\Delta)$ we denote by $\bar{v}$ the lift of $v$. This is a map defined on $C$ as follows, for all $(\lambda,x)\in\mathbb{R}_{>0}\times\Delta$, $\bar{v}(\lambda x)=\log(\lambda)+v(x)$. The map $v\mapsto\bar{v}$ is a bijection between $\operatorname{CLip}_{1}(\Delta)$ and $\mathscr{D}_{C}$. If we define $S$ on $\mathscr{D}_{C}$, then one verifies easily that, for all $(v,\lambda)\in\operatorname{CLip}_{1}(\Delta)\times\mathbb{R}$, $Fv\leqslant\lambda+v$ is equivalent to $S\bar{v}\leqslant\lambda+\bar{v}$ on $C$. So we can rewrite the dual characterization of the value of the game as follows.
Corollary 33.
The value $\rho$ of escape rate games played on cones has the following dual characterization
$\rho=\inf\{\lambda\in\mathbb{R}\mid\exists v\in\mathscr{D}_{C},Sv\leqslant\lambda+v\}$
In this way, we obtain the minimax characterization without the assumption on almost-isometries. However, there is no guarantee that there is a function $v\in\mathscr{D}_{C}$ verifying $Sv\leqslant\rho+v$, or $Fv\leqslant\rho+v$ —abusing notation and denoting by the same symbol $v$ a function on $\Delta$ and its lift. Nonetheless, if a continuous eigenfunction exists, then the associated eigenvalue is equal to the value of the game:
Proposition 34.
If
$Fv=\lambda+v$
is verified for some $\lambda\in\mathbb{R}$ and $v\in\mathscr{C}(\Delta)$, then $\lambda$ is necessarily equal to $\rho$.
Proof.
Indeed, in such a case the lifted function $v$ verifies $Sv=\lambda+v$ and we can construct stationary optimal strategies guaranteeing $\lambda$ in the game where the payoff is
$\limsup_{k\to\infty}\frac{v(T_{a_{k}b_{k}}\circ\dots\circ T_{a_{1}b_{1}}(x_{0}))}{k}$
by letting Player Min play $a_{k}\in\operatorname*{arg\,min}_{a\in\mathcal{A}}\max_{b\in\mathcal{B}}v(T_{ab}(x_{k-1}))$ and Player Max play $b_{k}\in\operatorname*{arg\,max}_{b\in\mathcal{B}}v(T_{a_{k}}b(x_{k-1}))$. Then as $v$ is continuous, it is bounded on $\Delta$ and because $v(\lambda x)=\lambda+v(x)$, for all $x\in\Delta$, we get that there exists $M\in\mathbb{R}^{+}$ such that for all $x\in X$
(18) $\log(\langle x,e^{*}\rangle)-M\leqslant v(x)\leqslant\log(\langle x,e^{*}\rangle)+M$
This inequality tells us this game is equivalent to $\Gamma_{\Delta}$ and therefore $\lambda=\rho$
Remark 35.
The eigenproblem above can be seen as a "2-player cohomological equation". In the classical cohomological equation [Livic1972], one seeks a function $f$ satisfying
$g=f\circ T-f$
where $g$ is a given function on a state space $X$ equipped with a transition map $T:X\to X$.
We next observe that an eigenfunction does exist under the following assumption.
Definition 36 (Tubular family).
Let $K\subset\operatorname{int}C$ be a cone such that the cross-section $\Delta_{K}\coloneqq\Delta\cap K$ is compact in the topology defined by the Thompson Metric. A family $(T_{ab})$ of self-maps of the interior of a cone $C$ that are nonexpansive in the Funk Metric is tubular with respect to $K$ if every map $T_{ab}$ preserves the cone $K$.
The terminology comes from the work of Bousch and Mairesse [Bousch_2001], who dealt with the special case in which $C$ is the standard orthant (up to a logarithmic change of variable). Gugliemi and Protasov [Guglielmi2011] and Jungers [Jungers2012] used the same notion, under the name of invariant embedded cone, to study the lower spectral radius.
In the next proposition, we denote by $\operatorname{CLip}_{1}(\Delta_{K})$ the set of functions defined on $\Delta_{K}$ that are $1$-Lipschitz with respect to the Funk metric induced by the cone $C$.
Proposition 37.
If the family $(T_{ab})_{(a,b)\in\mathcal{A}\times\mathcal{B}}$ is tubular with respect to the cone $K$, then there exists $v\in\operatorname{CLip}_{1}(\Delta_{K})$ such that
$Fv=\rho+v$
and the value of the escape rate game, $\rho$, is the unique eigenvalue of $F$.
Proof.
We consider the escape rate game with state space restricted to $K$ (including its boundary), equipped with the restriction of the Funk metric of the cone $C$ to the cone $K$. This game has the same value as the original conical game. We saw in 11 that there exist $\lambda\in\mathbb{R}$ and a function $v$ defined on $\operatorname{int}C$ that is $1$-Lipschitz with respect the Funk metric, and such that $Sv(x)=\lambda+v(x)$ holds for all $x\in\operatorname{int}C$. As $\Delta_{K}$ is compact, the function $v$ is bounded on $\Delta_{K}$, in particular, $m\coloneqq\min_{x\in\Delta_{K}}v(x)>-\infty$. By Lemma 28, this implies that $v(x)=v((\langle x,e^{*}\rangle)^{-1}x)+\log\langle x,e^{*}\rangle\geqslant\operatorname{Funk}(x,x_{0})-\beta+m$, and so $v$ is distance-like on $K$. Moreover, we have $[Fv](x)=\lambda+v(x)$ for all $x\in\Delta_{K}$. We get from 22 that $\lambda=\rho$ which proves the proposition. ∎
[…]
Proposition 38.
In the minimizer-free escape rate game the two following properties are equivalent
  1. (1)
    $\exists\alpha\in\mathbb{R}$ such that $\forall k\in\mathbb{N}^{*}$, $S^{k}d(\cdot,x_{0})\leqslant k\rho+d(\cdot,x_{0})+\alpha$
  2. (2)
    $\exists v\in\mathscr{D}$ such that $Sv\leqslant\rho+v$
Proof.
$(1)\implies(2)$. If (1) is verified we can define $v\coloneqq\sup_{k\in\mathbb{N}}S^{k}d(\cdot,x_{0})-k\rho$ which is distance-like as it is greater than $d(\cdot,x_{0})$. Since the game is assumed to be minimizer-free, the operator $S$ commutes with arbitrary suprema, and so
$Sv=\sup_{k\in\mathbb{N}}S^{k+1}d(\cdot,x_{0})-k\rho=\sup_{k\in\mathbb{N}^{*}}S^{k}d(\cdot,x_{0})-k\rho+\rho\leqslant v+\rho$
$(2)\implies(1)$. If (2) is verified, let $v\in\mathscr{D}$ such that $Sv\leqslant\rho+v$. Then there exists $\alpha\in\mathbb{R}$ such that for all $k\in\mathbb{N}*$
$S^{k}d(\cdot,x_{0})+\alpha\leqslant S^{k}v\leqslant k\rho+v\leqslant k\rho+d(\cdot,x_{0})+v(x_{0})\qed$
This result suggests that for escape rate games, the distance-like functions that achieve the infimum in 33 play the role of extremal norms.
6. The Example of Vector Addition Games
As an illustration, we give a complete solution of the vector addition game. We denote by $((\mathbb{R}^{n})^{*},\|\cdot\|_{*})$ the dual normed space of $(\mathbb{R}^{n},\|\cdot\|)$. Let us first notice that this game is trivial if $-\mathcal{B}\subset\mathcal{A}$. Then, for every $k\in\mathbb{N}^{*}$, Min can play tit-for-tat, setting $a_{k+1}\coloneqq-b_{k}$ at turn $k+1$, canceling the previous translation. This leads to a bounded trajectory, implying that the value of the escape rate is zero. We next denote by $\operatorname*{co}(\cdot)$ the convex hull of a set.
Proposition 39.
The vector addition game has a value equal to
(19) $\lambda\coloneqq\max_{b\in\mathcal{B}}\min_{a\in\operatorname*{co}(\mathcal{A})}\lVert a+b\rVert$
Any linear form $\ell$ achieving the maximum in
$\max_{\ell\in(\mathbb{R}^{n})^{*},\|\ell\|_{*}\leqslant 1}\max_{b\in\mathcal{B}}\min_{a\in\operatorname*{co}(\mathcal{A})}\ell(a+b)$
is a solution of the eigenproblem $S\ell=\lambda+\ell$. Furthermore, Max has a constant optimal strategy by playing
(20) $b^{*}\in\operatorname*{arg\,max}_{b\in\mathcal{B}}\min_{a\in\operatorname*{co}(\mathcal{A})}\lVert a+b\rVert$
Finally, the function
$\phi(x)=\text{dist}(x,-(n+1)\operatorname*{co}(\mathcal{A}))\enspace.$
is distance-like and verifies $S\phi\leqslant\lambda+\phi$. Hence, by playing at state $x\in\mathbb{R}^{n}$ an action $a$ which achieves the minimum in the expression of $S\phi(x)$, we obtain a stationary optimal strategy of Min.
Proof.
Using the characterization of the norm as a supremum over linear forms, together with Sion theorem on the commutation of min and max [Sion1958], we get:
$\lambda$ $=\max_{b\in\mathcal{B}}\min_{a\in\operatorname*{co}(\mathcal{A})}\lVert a+b\rVert=\max_{b\in\mathcal{B}}\min_{a\in\operatorname*{co}(\mathcal{A})}\max_{\ell\in(\mathbb{R}^{n})^{*},\|\ell\|_{*}\leqslant 1}\ell(a+b)$
$=\max_{b\in\mathcal{B}}\max_{\ell\in(\mathbb{R}^{n})^{*},\|\ell\|_{*}\leqslant 1}\min_{a\in\operatorname*{co}(\mathcal{A})}\ell(a+b)\ =\max_{\ell\in(\mathbb{R}^{n})^{*},\|\ell\|_{*}\leqslant 1}\max_{b\in\mathcal{B}}\min_{a\in\operatorname*{co}(\mathcal{A})}\ell(a+b)$
$=\max_{\ell\in(\mathbb{R}^{n})^{*},\|\ell\|_{*}\leqslant 1}\max_{b\in\mathcal{B}}\min_{a\in\mathcal{A}}\ell(a+b)=\max_{\ell\in(\mathbb{R}^{n})^{*},\|\ell\|_{*}\leqslant 1}(\max_{b\in\mathcal{B}}\ell(b)+\min_{a\in\mathcal{A}}\ell(a))$
$=\max_{\ell\in(\mathbb{R}^{n})^{*},\|\ell\|_{*}\leqslant 1}\min_{a\in\mathcal{A}}\max_{b\in\mathcal{B}}\ell(a+b)\enspace.$
Let $\ell$ be a linear form achieving the maximum in the expression above. Then, by definition of $S$, and using the linearity of $\ell$, we get
$S\ell$ $=\ell+\min_{a\in\mathcal{A}}\max_{b\in\mathcal{B}}\ell(a+b)=\ell+\lambda\enspace.$
Let Max play $b^{*}\in\operatorname*{arg\,max}_{b\in\mathcal{B}}\min_{a\in\operatorname*{co}(\mathcal{A})}\|a+b\|$, then for any $k\in\mathbb{N}^{*}$, we have that for all $(a_{1},\dots,a_{k})\in\mathcal{A}^{k}$
$\frac{\lVert x_{0}+a_{1}+b^{*}+\dots+a_{k}+b^{*}-x_{0}\rVert}{k}=\lVert\frac{a_{1}+\dots+a_{k}}{k}+b^{*}\rVert\geqslant\max_{b\in\mathcal{B}}\min_{a\in\operatorname*{co}(\mathcal{A})}\lVert a+b\rVert$
since $\frac{a_{1}+\dots a_{k}}{k}\in\operatorname*{co}(\mathcal{A})$
For the second inequality, let us first recall the useful Shapley Folkman lemma. Here, $\sum_{k=1}^{N}X_{k}$ denotes the Minkowski sum of subsets $X_{1},\dots,X_{N}\subset\mathbb{R}^{n}$.
Lemma 40 (Shapley Folkman [Starr1969]).
Let $N\in\mathbb{N}^{*}$ such that $N>n$. If $X_{1},\dots,X_{N}$ are nonempty bounded subset of $\mathbb{R}^{n}$ then for any $x\in\operatorname*{co}(\sum_{k=1}^{N}X_{k})=\sum_{k=1}^{N}\operatorname*{co}(X_{k})$, there exists $I\subset\{1,\dots,N\}$ such that $|I|\leqslant n$ and
$x=\sum_{i\in I}q_{i}+\sum_{k\in\{1,\dots,N\}/I}q_{k}$
where $q_{i}\in\operatorname*{co}(X_{i})/X_{i}$ and $q_{k}\in X_{k}$
Let $\tau\in\mathcal{B}^{s}$ a strategy of Max. To simplify the notations, we are going to denote $\max_{b\in\mathcal{B}}\min_{a\in\operatorname*{co}(\mathcal{A})}\lVert a+b\rVert=\max_{b\in\mathcal{B}}d(b,-\operatorname*{co}(\mathcal{A}))$ by $\lambda$. Given a strategy $\sigma$ of Min we will denote the resulting position after $k$ turns, i.e $x_{0}+a_{1}+b_{1}+\dots+a_{k}+b_{k}$ by $x_{k}(\sigma,\tau)$. We are going to prove that if $x_{0}\in-n\operatorname*{co}(\mathcal{A})$, there exists a strategy $\sigma^{*}\in\mathcal{A}^{s}$ such that for all $k\in\mathbb{N}$
$\delta_{k}\coloneqq d(x_{k}(\sigma^{*},\tau),-n\operatorname*{co}(\mathcal{A}))\leqslant k\lambda\enspace.$
This would mean that, starting from $-n\operatorname*{co}(\mathcal{A})$ we can guarantee a payoff lower than $\lambda$ and, as the initial position does not matter, we would have proven our proposition. Given the preceding sentence, we can consider the game where Max plays first.
Let us assume that for a certain $k\in\mathbb{N}$, the above inequality is true. It means that $x_{k}\in-n\operatorname*{co}(\mathcal{A})+kB(\lambda)$, where $B(\lambda)$ denotes the closed ball centered at 00 of radius $\lambda$, and since, by definition, the move $b_{k+1}$ played by Max given the strategy $\tau$ is in $-\operatorname*{co}(\mathcal{A})+B(\lambda)$. So
$x_{k}+b_{k+1}\in-(n+1)\operatorname*{co}(\mathcal{A})+(k+1)B(\lambda)$
The Shapley-Folkman lemma tells us that $-(n+1)\operatorname*{co}(\mathcal{A})\subset-n\operatorname*{co}(\mathcal{A})-\mathcal{A}$. So there exists at least one $\hat{a}\in\mathcal{A}$ such that
$x_{k}+b_{k+1}\in-n\operatorname*{co}(\mathcal{A})-\hat{a}+(k+1)B(\lambda)$
So by playing $a_{k}+1=\hat{a}$, we get that
$x_{k+1}\in-n\operatorname*{co}(\mathcal{A})+(k+1)B(\lambda)$
Then our proposition follows by induction. Now, this setting not only helps to better understand what escape rate games are, but it illustrates 21 and the fact that the distance-like function $\phi$ gives us the reverse inequality immediately. Indeed
$S\phi(x)$ $=\min_{a\in\mathcal{A}}\max_{b\in\mathcal{B}}d(x+a+b,-(n+1)\operatorname*{co}(\mathcal{A}))$
$=\min_{a\in\mathcal{A}}\max_{b\in\mathcal{B}}\min_{c\in(n+1)\operatorname*{co}(\mathcal{A})}\lVert x+a+b+c\rVert$
$\leqslant\min_{a\in\mathcal{A}}\min_{a^{\prime}\in n\operatorname*{co}(\mathcal{A})}\max_{b\in\mathcal{B}}\min_{d\in\operatorname*{co}(\mathcal{A})}\lVert x+a+a^{\prime}+b+d\rVert$
$\leqslant\min_{a\in\mathcal{A}}\min_{a^{\prime}\in n\operatorname*{co}(\mathcal{A})}\max_{b\in\mathcal{B}}\min_{d\in\operatorname*{co}(\mathcal{A})}\lVert x+a+a^{\prime}\rVert+\lVert b+d\rVert$
[…]
$=\phi(x)+\max_{b\in\mathcal{B}}\min_{d\in\operatorname*{co}(\mathcal{A})}\lVert b+d\rVert=\phi(x)+\lambda\enspace.\qed$
7. Concluding remarks
We finally raise some open questions. We observed that when the non-linear eigenproblem $Sv=\lambda+v$ is solvable with a function $v$ that is distance like, then, $\lambda$ is unique and coincides with the value of the escape rate game. It would interesting to give conditions for the solvability of this eigenproblem, beyond the tubular setting of Proposition 37; we note that related issues have arisen in the study of the cohomological equation and its generalizations, in the context of dynamical systems [Livic1972, Bousch2001, Bousch2011].
Whereas, in the escape rate game, Player Max always has an optimal stationary stragegy, we only showed that Player Min has an optimal stationary strategy when the non-linear eigenproblem $Sv\leqslant\rho+v$ is solvable for some distance-like function $v$. Investigating the existence of an optimal stationary strategy of Player Min beyond this case remains an open question.
[…]